### v5

# Seifert surfaces in the four-ball and composition of binary quadratic forms

Abstract.
We use composition of binary quadratic forms to systematically create pairs of Seifert surfaces that are non-isotopic in the four-ball. Our main topological result employs Gauss composition to classify the pairs of binary quadratic forms that arise as the Seifert forms of pairs of disjoint Seifert surfaces of genus one. The main ingredient of the proof is number-theoretic and of independent interest. It establishes a new connection between the Bhargava cube and the geometric approach to Gauss composition via planes in the space of two-by-two matrices. In particular, we obtain a geometric recipe that given any two binary quadratic forms finds a Bhargava cube that gives rise to their composition.
Key words and phrases:
Bhargava Cube, Gauss composition, Klein vectors, Seifert surfaces, Seifert form
2020 Mathematics Subject Classification:
57K10, 57K45, 11E16, 15A72
PF was supported by the SNSF grant 181199. AW was funded by ERC grant HomDyn, ID 833423.
1. Introduction
In this article, we revisit Gauss composition of binary quadratic forms via planes in four-dimensional space motivated by applications which obtain robust constructions of Seifert surfaces that are non-isotopic in the four-ball. We begin here by discussing these applications in §1.1 but note that the number-theoretic result in §1.2 is of independent interest. For example, the latter leads to a curious variant of the mixing conjecture of Michel and Venkatesh [22]. In fact, the number-theoretic parts of the paper (§1.2, §1.3 and §1.4 in the introduction, and §2, §3, §4, §7.1 in the body of the text) can be read completely independently of the low-dimensional topology part of the paper. The latter is written to be independent but makes reference to some notation and results of the number-theoretic parts.
1.1. Seifert surfaces with the same boundary
Recently, Hayden, Miller, Kim, Park, and Sundberg [17] found a pair of genus one Seifert surfaces $F_{1}$ and $F_{2}$ for the same knot $K$ that, after pushing their interior into the four-ball $B^{4}$, are not topologically ambient isotopic in $B^{4}$. Here, recall that a Seifert surface for a knot $K$ is a compact oriented smooth surface in the $3$-sphere $S^{3}$ with boundary $K$. This resolved a question by Livingston [21], who asked whether such two Seifert surfaces can exist. The example in [17] is such that the interiors are disjoint; we call pairs of genus 1 Seifert surfaces with the same boundary and disjoint interior disjoint genus one pairs from here on. In this paper, we put the recent example into a more general context by producing algebraic characterizations of when such pairs of Seifert surfaces with prescribed classical invariants (derived from the Seifert form) exist. The key result that makes all our topology results possible is a complete number-theoretic characterization of those pairs of Seifert forms that arise from disjoint genus one pairs of Seifert surfaces.
It turns out that this characterization is in terms of Gauss composition for integral binary quadratic forms [15], which we denote by *. To state our characterization, we recall that the discriminant of an integral binary quadratic form (IBQF) $ax^{2}+bxy+cy^{2}$ is $b^{2}-4ac$. Furthermore, the Seifert form is a bilinear form associated with a Seifert surface. In the case of a genus one Seifert surface of a knot it amounts to an $\operatorname{SL}_{2}(\mathbb{Z})$-equivalence class $ax^{2}+bxy+cy^{2}]$ of an IBQF; compare (5.1).
Theorem 1.1.
Let $s_{1}$ and $s_{2}$ be $\operatorname{SL}_{2}(\mathbb{Z})$-equivalence classes of IBQFs. The pair $s_{1}$ and $s_{2}$ arises as the pair of Seifert forms of a disjoint genus one pair of Seifert surfaces if and only if there exist $a,c\in\mathbb{Z}$ such that $s_{1}$ and $s_{2}$ have discriminant 14ac1-4ac and $ax^{2}+xy+cy^{2}]^{2}*s_{1}=s_{2}$.
Before we explain a purely algebraic formulation of Theorem 1.1 and how Gauss composition, in fact Bhargava’s cube description of it, arises, we describe consequences concerning non-isotopy of pushed-in Seifert surfaces. For this we use the if-statement of Theorem 1.1.
Arguably the simplest classical integer-valued knot invariant is the determinant, denoted by $D_{K}$ for a knot $K$. We choose the sign convention for which $D_{K}\equiv 1\,\mathrm{mod}\,{4}$ for all knots $K$. With this convention, if a genus one Seifert surface for $K$ has Seifert form $ax^{2}+bxy+cy^{2}]$, then $D_{K}$ equals the discriminant of $ax^{2}+bxy+cy^{2}$. Recall that for a fixed non-zero integer $D$, the oriented class group $\mathcal{G}^{+}_{D}$ is the set of the $\operatorname{SL}_{2}(\mathbb{Z})$-equivalence classes of primitive IBQFs—those with coprime coefficients—with discriminant $D$, which form an abelian group under Gauss composition. Note that if $D$ is square-free, every IBQF with discriminant $D$ is primitive. For $D\equiv 1\,\mathrm{mod}\,{4}$, the neutral element is $1=[x^{2}+xy+\frac{-D+1}{4}y^{2}]$.
Theorem 1.2.
Fix an integer $D\equiv 1\,\mathrm{mod}\,{4}$ and suppose that there exists $a,c\in\mathbb{Z}$ with 14ac=D1-4ac=D such that $ax^{2}+xy+cy^{2}]^{2}\neq 1$. Then there exists a knot $K$ with determinant $D$ that has two disjoint genus 1 Seifert surfaces that are not ambiently isotopic in $B^{4}$ after being pushed in the $4$-ball.
Suppose that in addition there exists $a^{\prime},c^{\prime}\in\mathbb{Z}$ with $1-4a^{\prime}c^{\prime}=D$ such that
$ax^{2}+xy+cy^{2}]^{2}\neq[a^{\prime}x^{2}+xy+c^{\prime}y^{2}]^{2}\neq 1.$
Then one of the above Seifert surfaces can be chosen such that its Seifert form is any given class of a primitive IBQF of discriminant $D$.
Theorem 1.2 should be seen as a special case of the robust construction given in this article. In particular, we give an exact characterization for when one can distinguish disjoint pairs of genus one Seifert surfaces by the intersection form of the double branched cover. Also, the primitivity assumption in the theorem is removed below using more language concerning IBQFs which we avoid here for sake of exposition (cf. Theorem 5.7).
It turns out the condition $ax^{2}+xy+cy^{2}]^{2}\neq 1$ is often easy to check in the oriented class group. In fact, we have the following explicit result for negative determinants. (We note that given our sign convention, negative determinants corresponds to the case of a definite symmetrized Seifert from or, equivalently, knots with non-zero signature.)
Corollary 1.3.
Fix a negative integer $D\equiv 1\,\mathrm{mod}\,{4}$. There exist a knot $K$ with determinant $D$ and two Seifert surfaces with boundary $K$ that are not ambiently isotopic in $B^{4}$ after being pushed in the $4$-ball if
$\frac{-D+1}{4}\notin\left\{p^{k}\mid p\text{ prime and }k\in\{0,1,2\}\right\}.$
Example 1.4.
The largest negative integer $D$ with $D\equiv 1\,\mathrm{mod}\,{4}$ such that $\frac{-D+1}{4}$ is not $1$, a prime, or a square of a prime is 23-23. The example in [17] corresponds exactly to this case. Up to taking inverses there is exactly one non-trivial element in $\mathcal{G}^{+}_{-23}$ that is the square of an element of the form $ax^{2}+xy+cy^{2}]$:
$2x^{2}\pm xy+3y^{2}]^{2}=[-(2x^{2}\pm xy+3y^{2})]^{2}=[2x^{2}\mp xy+3y^{2}]\neq 1=[x^{2}+xy+6y^{2}].$
Hence, by Theorem 1.2, there must exist pairs of Seifert surfaces for a knot with determinant 23-23 that are not isotopic in $B^{4}$ after pushing in, as is claimed in Corollary 1.3. See Example 6.2 for more details and the explicit connection to the example from [17].
Fehér [13] has exhibited more examples of disjoint genus one pairs of Seifert surfaces. We put these examples in the number-theoretic framework of this article (cf. Example 6.4).
Our framework readily allows treating positive discriminants as well.
Example 1.5.
The smallest positive discriminant $D\equiv 1\,\mathrm{mod}\,4$ for which Theorem 1.2 provides an example is $D=25$. There exists an element of $\mathcal{G}_{25}^{+}$ which is a non-trivial square of an element of the form $ax^{2}+xy+cy^{2}]$, namely $x^{2}+5xy]$ (cf. §6.3). Hence, Theorem 1.2 applies. In fact, the above can be extended to all odd square discriminants $D\geq 25$; see Corollary 6.10.
The key algebraic object that governs which Seifert forms arise from disjoint genus one pairs is a particular bilinear form on $\mathbb{Z}^{4}$. This bilinear form is in fact the Seifert form of a closed genus 2 surface in $S^{3}$. Below we describe it in terms of its symmetrization $Q$ and antisymmetrization $\theta$ and all that follows are results about this bilinear form that are independent of the knot theory above.
1.2. Gauss composition via planes in $\mathbb{Z}^{4}$
Consider $\mathbb{Z}^{4}$ equipped with the quadratic form
$Q(x_{1},x_{2},x_{3},x_{4})=x_{1}x_{2}+x_{3}x_{4}.$
We call a rank two oriented summand $L\subset\mathbb{Z}^{4}$ a plane.
Following [1], we parametrize planes $L$ by two so-called Klein vectors $a_{1}(L),a_{2}(L)$. In our setting, the latter can be thought of as a pair of IBQFs of equal discriminant. The correspondence between these pairs of IBQFs and planes is given by explicit formulas both ways (cf. §3). For an equivalence class $[q]$ we denote by $\overline{[q]}=[q(-x,y)]$ the orientation-reversed class. The following is the main number-theoretic result in this article.
Theorem 1.6.
For any plane $L$ with $\operatorname{disc}(Q|_{L})\neq 0$, we have
(1.1) $[Q|_{L}]=\overline{[a_{1}(L)]}*[a_{2}(L)].$
One may see the above theorem as a geometric point of view on Gauss composition. Indeed, here is a concrete realization of Gauss composition: Given a pair of, say, primitive IBQFs $q_{1}=a_{1}x^{2}+b_{1}xy+c_{1}y^{2},q_{2}=a_{2}x^{2}+b_{2}xy+c_{2}y^{2}$ of equal discriminant, consider
(1.2) $L=\left\{M\in\operatorname{M}_{2}(\mathbb{Z}):\big(\begin{smallmatrix}-b_{1}&-2a_{1}\\ 2c_{1}&b_{1}\end{smallmatrix}\big)M=M\big(\begin{smallmatrix}b_{2}&-2a_{2}\\ 2c_{2}&-b_{2}\end{smallmatrix}\big)\right\}$
where $\operatorname{M}_{2}$ denotes the algebra of two by two matrices. Since $q_{1}$ and $q_{2}$ have the same discriminant, $L$ is a free rank two module over $\mathbb{Z}$, which is given as the solution of the four concrete linear equations in (1.2). One finds a basis $v_{1},v_{2}$ of $L$. Up to switching these basis vectors, the composition of $q_{1}]$ and $q_{2}]$ is $q_{1}]*[q_{2}]=[q]$ where $q(x,y)=\det(xv_{1}+yv_{2})$; see §4 for more details.
To summarize, composing two classes amounts to computing a basis of an explicit plane. We remark that classical forms of Gauss composition usually reduce to solving congruence equations modulo the first coefficients of $q_{1},q_{2}$ (which can be arranged to be coprime). Upon close consideration, a similar statement is true for the above basis search after elementary manipulations.
We provide two proofs of Theorem 1.6; one giving a new geometric perspective on Dirichlet composition and the other using Bhargava’s cube [3]. In fact, the latter gives a geometric recipe that given any two IBQFs finds a Bhargava cube [3] that gives rise to their composition (cf. Remark 4.10).
Remark 1.7.
Let us briefly remark on the historical context surrounding Theorem 1.6. In [1], planes and their associated integral quadratic forms were studied from a viewpoint of equidistribution (typical statistical independence of $Q_{L}]$ and $Q_{L^{\perp}}]$, where $L^{\perp}$ is the orthogonal complement with respect to $Q$). Here, the ambient quadratic form $Q$ can be any norm form on a quaternion algebra over $\mathbb{Q}$. In the process of the proof, the authors found a version of the relations in Theorem 1.6 (see [1, Thm. 8.1& App.B]) though the exact formulation was irrelevant to the equidistribution purpose. Chinta and Pereira Júnior [8] later proved a slightly weaker form of these relations for the quadratic form $Q=x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{4}^{2}$. The proof given here of the relations in (1.1) could be carried out via the adelic machinery established in [1].
1.3. Symplectic planes and applications to topology
In addition to the quadratic form $Q$, we equip $\mathbb{Z}^{4}$ with the integral symplectic form $\theta(\cdot,\cdot)$ determined by
$\theta(e_{1},e_{2})=-\theta(e_{2},e_{1})=\theta(e_{3},e_{4})=-\theta(e_{4},e_{3})=1,$
$\theta(e_{i},e_{j})=0\text{ for }\{i,j\}\notin\left\{\{1,2\},\{3,4\}\right\}.$
where $(e_{1},e_{2},e_{3},e_{4})$ is the standard basis. A plane $L$ is called symplectic if it has an oriented basis $(v_{1},v_{2})$ with $\theta(v_{1},v_{2})=1$. Its symplectic complement is defined as $L^{\perp\kern-4.2pt\perp}=\left\{x\in\mathbb{Z}^{4}:\theta(x,y)=0\text{ for all }y\in L\right\}$ (with the orientation that makes it symplectic). As we will see in §5, in order to establish the existence of a disjoint genus one pair of Seifert surface for a knot with prescribed determinant, it suffices to answer the following question. For which discriminants is there a symplectic plane $L$ with
(1.3) $[Q|_{L}]\neq[Q|_{L^{\perp\kern-4.2pt\perp}}],\overline{[Q|_{L^{\perp\kern-4.2pt\perp}}]}\,?$
In fact, we prove in particular the following.
Corollary 1.8.
Given an integer $D\equiv 1\,\mathrm{mod}\,4$, there exists a symplectic plane $L\subset\mathbb{Z}^{4}$ such that $Q|_{L}$ has discriminant $D$ and $Q|_{L}]\neq[Q|_{L^{\perp\kern-4.2pt\perp}}],[\overline{Q|_{L^{\perp\kern-4.2pt\perp}}}]$ if and only if there exists $a,c\in\mathbb{Z}$ with 14ac=D1-4ac=D and $ax^{2}+xy+cy^{2}]^{2}\neq 1$.
Corollary 1.8 and variants thereof are the main input needed for the theorems in §1.1 as we will see in §5. Let us briefly discuss how (1.3) relates to Theorem 1.6.
Sketch of proof of Corollary 1.8.
For a plane $L$, a simple observation (cf. §3.3) shows that $a_{2}(L)$ is an IBQF of the form $ax^{2}+xy+cy^{2}$ if and only if $L$ is symplectic. Moreover, in this case $a_{1}(L^{\perp\kern-4.2pt\perp})=-a_{1}(L)$ and $a_{2}(L^{\perp\kern-4.2pt\perp})=-\overline{a_{2}(L)}$ as can be shown by elementary calculations (cf. Proposition 3.6). Altogether, applying Theorem 1.6 twice (for $L$ and $L^{\perp\kern-4.2pt\perp}$) we get
(1.4) $[a_{2}(L)]^{-2}*[Q|_{L}]=\overline{[Q|_{L^{\perp\kern-4.2pt\perp}}].}$
Choosing $a_{1}(L)]=[a_{2}(L)]$ one obtains a symplectic plane as in (1.3) whenever there exists an IBQF of the form $ax^{2}+xy+cy^{2}$ of the same discriminant with $ax^{2}+xy+cy^{2}]^{2}\neq 1$. ∎
1.4. The mixing conjecture and symplectic planes
Consider the collection of pairs
(1.5) $\big\{([Q|_{L}],[Q|_{L^{\perp\kern-4.2pt\perp}}]):L\text{ symplectic},\,\operatorname{disc}(Q|_{L})=D\big\}.$
As we show in §5, these are (up to an orientation-change in the second component) the pairs of Seifert forms of a disjoint genus one pair of Seifert surfaces as considered in §1.1. Moreover, as is visible from the above sketch of proof of Corollary 1.8, the set (1.5) is a very sparse subset of the set of all pairs of classes of IBQF’s of discriminant $D$. This sparsity gives rise to an equidistribution problem fitting into the context of the mixing conjecture posed by Michel and Venkatesh [22], which we phrase in this subsection.
Any $\operatorname{SL}_{2}(\mathbb{Z})$-class $s$ of a positive-definite IBQF gives rise to a complex multiplication (CM) point on the modular curve
$\mathrm{z}_{s}\in Y_{0}(1)=\operatorname{SL}_{2}(\mathbb{Z})\backslash\mathbb{H}$
Explicitly, the CM point associated to $ax^{2}+bxy+cy^{2}]$ is the $\operatorname{SL}_{2}(\mathbb{Z})$-orbit under Möbius transformations of the unique root of $ax^{2}+bx+c$ contained in the complex upper-half plane $\mathbb{H}$. We wish to examine for $D<0$ with $D\equiv 1\,\mathrm{mod}\,4$ the distribution of the set of pairs
(1.6) $\mathcal{CM}_{D,\mathrm{sympl}}=\big\{(\mathrm{z}_{[Q|_{L}]},\mathrm{z}_{[Q|_{L^{\perp\kern-4.2pt\perp}}]}):L\text{ symplectic},\,\operatorname{disc}(Q|_{L})=D\big\}$
in the product $Y_{0}(1)\times Y_{0}(1)$. Let
$\mathcal{C}_{D}=\{[ax^{2}+xy+cy^{2}]^{2}:a,c\in\mathbb{Z}\text{ with }1-4ac=D\}.$
Conjecture 1.9.
Let $(D_{i})_{i}$ be a divergent sequence of negative discriminants with $D_{i}\equiv 1\,\mathrm{mod}\,4$. Suppose additionally that $|\mathcal{C}_{D_{i}}|\to\infty$. Then the finite sets $\mathcal{CM}_{D_{i},\mathrm{sympl}}$ are equidistributed in $Y_{0}(1)\times Y_{0}(1)$ with respect to the product of the normalized hyperbolic measures. That is, for any $f\in C_{c}(Y_{0}(1)\times Y_{0}(1))$ we have as $i\to\infty$
$\frac{1}{|\mathcal{CM}_{D_{i},\mathrm{sympl}}|}\sum_{(\mathrm{z}_{1},\mathrm{z}_{2})\in\mathcal{CM}_{D_{i},\mathrm{sympl}}}f(\mathrm{z}_{1},\mathrm{z}_{2})\to\int_{Y_{0}(1)\times Y_{0}(1)}f.$
It follows from (1.4) that
(1.7) $\mathcal{CM}_{D,\mathrm{sympl}}=\{(\mathrm{z}_{s},\mathrm{z}_{c\ast\overline{s}}):c\in\mathcal{C}_{D},\ s\text{ a $\operatorname{SL}_{2}(\mathbb{Z})$-class of IBQFs of discriminant }D\}.$
Thus, the condition $|\mathcal{C}_{D_{i}}|\to\infty$ in Conjecture 1.9 equivalently asserts that the proportion of points in $\mathcal{CM}_{D_{i},\mathrm{sympl}}$ contained in any fixed Hecke correspondence goes to zero as $i\to\infty$. In particular, the condition is necessary. Also, notice that equidistribution in each of the factors $Y_{0}(1)$ in Conjecture 1.9 follows from a theorem of Duke [9].
By (1.7), Conjecture 1.9 is a special case of the mixing conjecture of Michel and Venkatesh [22] which has seen dramatic progress in recent years starting with Khayutin’s breakthrough [19] using fundamental rigidity results for diagonalizable actions on locally homogeneous spaces due to Einsiedler and Lindenstrauss [10]. Subsequent work of Blomer and Brumley [5], Blomer, Brumley, and Khayutin [4], and the upcoming work of Blomer, Brumley, Radziwiłł employed purely analytic methods to deduce under increasingly weaker assumptions on zeros of certain $L$-functions. The question arises whether the additional average in Conjecture 1.9 (namely over $\mathcal{C}_{D_{i}}$) makes the problem more accessible. For instance, one could aim to prove Conjecture 1.9 assuming no Landau-Siegel zeros for Dirichlet $L$-functions associated to the Kronecker symbols of the fields $\mathbb{Q}(\sqrt{D_{i}})$.
Organization of the paper
This paper is organized as follows.
In §2 we discuss some notation valid throughout the article.
In §3, we reconstruct the Klein map from [1, §2] for planes in the algebra of $2\times 2$ matrices. We find the restriction on values of the Klein map at symplectic planes in §3.3.
In §4, we establish relations between the quadratic forms of planes and their associated Klein vectors. In particular, this permits a comparison between the quadratic forms of planes and their orthogonal (or symplectic) complements.
In §5, we recall various notions from knot theory and prove Theorem 1.1 and 1.2. We also discuss our results in the context of S-equivalence (in §5.4) and the previous results of the third author [23] as well as the realization problem for a fixed knot (in §5.5).
In §6, we provide various examples to illustrate our main theorems for varying discriminants. In particular, we prove Corollary 1.3 in §6.1.
In §7, we discuss a variety of open problems relating to the results of the present article.
Acknowledgments
We thank Ken Baker, Gautam Chinta, Manfred Einsiedler, Zsombor Fehér, Kyle Hayden, and Filippos Sytilidis for helpful discussions. We in particular thank Zsombor for his very careful reading of an earlier version and for detailed feedback.
2. Notation
We collect some notation, pertaining mostly to integral binary quadratic forms (IBQFs). For an IBQF $q(x,y)=ax^{2}+bxy+cy^{2}$ we recall that the discriminant is $\operatorname{disc}(q)=b^{2}-4ac=-4\det(M)$ when $M$ is the symmetric matrix representing $q$. An integer $D$ is correspondingly called a discriminant if $D\equiv 0,1\,\mathrm{mod}\,4$. In the topological part of this article, one may safely assume $D\equiv 1\,\mathrm{mod}\,4$. The content of $q$, denoted by $\operatorname{content}(q)\in\mathbb{Z}_{\geq 1}$, is defined to be the greatest common divisor of the coefficients of $q$. We say that $q$ is primitive if its content is $1$.
The group $\operatorname{GL}_{2}(\mathbb{Z})$ acts by composition on the set of IBQFs preserving the discriminant as well as the content. The $\operatorname{SL}_{2}(\mathbb{Z})$-equivalence class of an IBQF $q$ is denoted by $[q]$. We denote by $\mathcal{Q}_{D}^{+}$ the set of $\operatorname{SL}_{2}(\mathbb{Z})$-equivalence classes of IBQFs of discriminant $D$ and by $\mathcal{G}_{D}^{+}$ the subset of equivalence classes of primitive forms. Gauss composition defines a composition rule on $\mathcal{G}_{D}^{+}$ which turns it into an abelian group. Note that it does not yield a composition rule on $\operatorname{GL}_{2}(\mathbb{Z})$-equivalence classes. We refer to $\mathcal{G}_{D}^{+}$ equipped with Gauss composition as the oriented class group of discriminant $D$ henceforth. We write $\bar{q}(x,y)=q(-x,y)$ for a choice of orientation reversal and recall that $\bar{q}]=[q]^{-1}$ for $q]\in\mathcal{G}_{D}^{+}$. Also recall that for $D\equiv 1\,\mathrm{mod}\,4$ (resp. $D\equiv 0\,\mathrm{mod}\,4$) the identity element of $\mathcal{G}_{D}^{+}$ is $1=1_{D}=[x^{2}+xy-\frac{D-1}{4}y^{2}]$ (resp. $1=1_{D}=[x^{2}-\frac{D}{4}y^{2}]$).
Remark 2.1.
The oriented class group $\mathcal{G}_{D}^{+}$ factors onto the usual class group $\mathcal{G}_{D}$. In fact, $\mathcal{G}_{D}$ can be seen as the quotient of the set of primitive IBQFs of discriminant $D$ by the twisted $\operatorname{GL}_{2}(\mathbb{Z})$-action through $g.q=\frac{1}{\det(g)}q\circ g^{t}$. Alternatively, $\mathcal{G_{D}}$ is the quotient of $\mathcal{G}_{D}^{+}$ by the $\mathbb{Z}/2\mathbb{Z}$-action given by $q]\mapsto[-\overline{q}]$. In particular, the map $\mathcal{G}_{D}^{+}\to\mathcal{G}_{D}$ is at most 212-1 and we have the following.
  • When $D<0$, the $\operatorname{SL}_{2}(\mathbb{Z})$-action preserves the signature and hence the map $\mathcal{G}_{D}^{+}\to\mathcal{G}_{D}$ is 212-1.
  • When $D>0$ and a solution to the negative Pell equation exists, $\mathcal{G}_{D}^{+}\to\mathcal{G}_{D}$ is an isomorphism. Recall that for $D\equiv 1\,\mathrm{mod}\,4$ (resp. $D\equiv 0\,\mathrm{mod}\,4$) the negative Pell equation is $x^{2}+xy-\frac{D-1}{4}y^{2}=-1$ (resp. $x^{2}-\frac{D}{4}y^{2}=-1$). If no solution exists, it is again 212-1. In terms of hyperbolic geometry, this criterion amounts to saying that the geodesic associated to $q$ (obtained through connecting the zeros of $q$ on the boundary $\mathbb{P}^{1}(\mathbb{R})\simeq\partial\mathbb{H}$) is equivalent via $\operatorname{SL}_{2}(\mathbb{Z})$ to its time reversal.
We also remark that Gauss composition can be extended to a composition rule for IBQFs of coprime content (cf. §4.2).
For $D\equiv 1\,\mathrm{mod}\,4$, define $\mathcal{S}^{+}_{D}$ to be the subgroup of $\mathcal{G}^{+}_{D}$ generated by the set
(2.1) $\{[ax^{2}+xy+cy^{2}]^{2}:a,c\in\mathbb{Z}\text{ with }1-4ac=D\}.$
Note that (2.1) (and hence also $\mathcal{S}^{+}_{D}$) is invariant under $q]\mapsto[\bar{q}]$ as $ax^{2}-xy+cy^{2}]=[cx^{2}+xy+ay^{2}]$ for any $a,c$.
Binary forms may be identified with traceless matrices as follows. Denote by $\operatorname{M}_{2}$ the space of $2\times 2$ matrices and by $\operatorname{M}_{2}^{0}$ the subspace of traceless matrices in $\operatorname{M}_{2}$. Define the so-called Gross lattice
$\Lambda=\left\{a\colon a\in\big(\begin{smallmatrix}\mathbb{Z}&2\mathbb{Z}\\ 2\mathbb{Z}&\mathbb{Z}\end{smallmatrix}\big)\cap\operatorname{M}_{2}^{0}\right\}\subset\operatorname{M}_{2}^{0}(\mathbb{Z}).$
There is a linear bijection between the set of IBQFs and $\Lambda$. Indeed, we identify a binary quadratic form $q=ax^{2}+bxy+cy^{2}$ with the traceless matrix $A(q)=\big(\begin{smallmatrix}b&-2a\\ 2c&-b\end{smallmatrix}\big)$. We denote the inverse of this identification by $A\in\operatorname{M}_{2}^{0}\mapsto q_{A}$. The identification is equivariant with respect to the $\operatorname{SL}_{2}(\mathbb{Z})$-actions where one acts by conjugation on $\operatorname{M}_{2}^{0}(\mathbb{Z})$, that is, $A(g.q)=gA(q)g^{-1}$. Note that $A(\cdot)$ restricts to an equivariant map from $\{q:\operatorname{disc}(q)=D\}$ to traceless matrices in $\Lambda$ of determinant D-D. Moreover, $A(q)$ is primitive in the Gross lattice $\Lambda$ if and only if $q$ is primitive.
3. The Klein map
In this section, we set up the correspondence between planes in $\mathbb{Z}^{4}$ and Klein vectors as in [1]. This was already alluded to in the introduction. In §3.3, we specialize the construction to symplectic planes.
3.1. Setup and Notation
We may identify $\mathbb{R}^{4}$ with the space $\operatorname{M}_{2}(\mathbb{R})$ and $\mathbb{Z}^{4}$ with $\operatorname{M}_{2}(\mathbb{Z})$. The usual (standard) involution on $\operatorname{M}_{2}$ is the adjugate operation, which we denote by $\overline{\big(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\big)}=\big(\begin{smallmatrix}d&-b\\ -c&a\end{smallmatrix}\big)$. With respect to the basis $\mathcal{B}=\left(\big(\begin{smallmatrix}1&0\\ 0&0\end{smallmatrix}\big),\big(\begin{smallmatrix}0&0\\ 0&1\end{smallmatrix}\big),\big(\begin{smallmatrix}0&0\\ -1&0\end{smallmatrix}\big),\big(\begin{smallmatrix}0&1\\ 0&0\end{smallmatrix}\big)\right)$ the bilinear form
$M(x,y)=\operatorname{Tr}\left(x\big(\begin{smallmatrix}1&0\\ 0&0\end{smallmatrix}\big)\overline{y}\right)$
is given by the matrix
(3.1) $U=\begin{pmatrix}0&1&0&0\\ 0&0&0&0\\ 0&0&0&1\\ 0&0&0&0\end{pmatrix}$.
Since $M(y,x)=\operatorname{Tr}\left(y\big(\begin{smallmatrix}1&0\\ 0&0\end{smallmatrix}\big)\overline{x}\right)=\operatorname{Tr}\left(\overline{y\big(\begin{smallmatrix}1&0\\ 0&0\end{smallmatrix}\big)\overline{x}}\right)=\operatorname{Tr}\left(x\big(\begin{smallmatrix}0&0\\ 0&1\end{smallmatrix}\big)\overline{y}\right)$ the symmetric form $Q=M+M^{t}$ and the symplectic form $\theta=M-M^{t}$ are given by
(3.2) $Q(x,y)=\operatorname{Tr}\left(x\overline{y}\right),\quad\theta(x,y)=\operatorname{Tr}\left(xj\overline{y}\right).$
where $j$ denotes the matrix $\big(\begin{smallmatrix}1&0\\ 0&-1\end{smallmatrix}\big)$. As customary for quadratic forms, we set $Q(x)=Q(x,x)$ and note that $Q(x)=2\det x$.
3.2. Planes and the Klein map
Recall that a plane $L\subset\mathbb{Z}^{4}$ is a rank two oriented summand. Here, an orientation amounts to a choice of basis of $L$ considered up to the equivalence relation of applying integral basis changes with determinant 1. In the following, any basis chosen for a plane will be taken to match the orientation. Via the identification $\mathbb{Z}^{4}\simeq\operatorname{M}_{2}(\mathbb{Z})$ from above, any plane will also be considered an oriented summand of $\operatorname{M}_{2}(\mathbb{Z})$. A plane $L$ defines an equivalence class $q_{L}]$, where $q_{L}$ is defined via a choice of basis $(v_{1},v_{2})$ of $L$ as follows
(3.3) $q_{L}(x,y):=\frac{1}{2}Q(xv_{1}+yv_{2})=\frac{1}{2}\left(Q(v_{1})x^{2}+2Q(v_{1},v_{2})xy+Q(v_{2})y^{2}\right)=\\ =\det(v_{1})x^{2}+\operatorname{Tr}(v_{1}\overline{v_{2}})xy+\det(v_{2})y^{2}\in\mathbb{Z}[x,y]$
Let $\mathcal{L}$ denote the set of planes. We set $\mathcal{L}_{\neq 0}=\{L\in\mathcal{L}:\operatorname{disc}(q_{L})\neq 0\}$ and $\mathcal{L}_{D}=\left\{L\in\mathcal{L}:\operatorname{disc}(q_{L})=D\right\}$ for every non-zero discriminant $D$. We note that $\mathcal{L}_{D}$ is always non-empty; e.g. for $\equiv 1\,\mathrm{mod}\,4, D\mathcal{L}_{D}$ contains the plane spanned by $\big(\begin{smallmatrix}1&0\\ 0&1\end{smallmatrix}\big)$ and $\big(\begin{smallmatrix}1&\frac{D-1}{4}\\ 1&0\end{smallmatrix}\big)$. Similarly to [1] we would like to study $\mathcal{L}_{D}$ via two associated vectors in $\operatorname{M}_{2}^{0}(\mathbb{Z})$, which we call the associated Klein vectors. The reader can compare this discussion to [1, §2] but notice that here we consider subspaces (or sublattices) with orientation (cf. also [1, §7]). Note that $\operatorname{SL}_{2}(\mathbb{Z})\times\operatorname{SL}_{2}(\mathbb{Z})$ acts on $\operatorname{M}_{2}(\mathbb{Z})$ by $(g_{1},g_{2}).x=g_{1}xg_{2}^{-1}$ which clearly preserves the determinant. In particular, this action induces an action of $\operatorname{SL}_{2}(\mathbb{Z})\times\operatorname{SL}_{2}(\mathbb{Z})$ on $\mathcal{L}_{D}$ for every $D$.
Given a plane $L$ we write $L_{\mathrm{op}}$ for the same plane with reversed orientation. Note $q_{L_{\mathrm{op}}}]=[\overline{q_{L}}]$. The orthogonal complement $L^{\perp}$ of a plane $L$ is defined as
(3.4) $L^{\perp}=\left\{x\in\operatorname{M}_{2}(\mathbb{Z})\colon Q(x,\ell)=0,\,\forall\ell\in L\right\}$
with orientation as declared in (3.7) below.
For a plane $L\in\mathcal{L}$ we choose a basis $(v_{1},v_{2})$ for $L$ and consider the traceless matrices
(3.5) $a_{1}(L)=2v_{1}\overline{v_{2}}-\operatorname{Tr}\left(v_{1}\overline{v_{2}}\right)I_{2},\quad a_{2}(L)=2\overline{v_{2}}v_{1}-\operatorname{Tr}\left(\overline{v_{2}}v_{1}\right)I_{2}.$
We call $a_{1}(L)$ and $a_{2}(L)$ the Klein vectors (“vectors” in the three-dimensional vector space $\operatorname{M}_{2}^{0}$) associated to $L$. By definition, $a_{1}(L)$ and $a_{2}(L)$ are elements of the Gross lattice $\Lambda$.
Definition 3.1 (pair-primitivity).
We call two matrices $a_{1},a_{2}\in\Lambda$ pair-primitive if $(a_{1},a_{2})$ is a primitive element of $\Lambda^{2}\subset\operatorname{M}_{2}^{0}(\mathbb{Q})^{2}$. That is, there is no prime $p$ such that $\tfrac{1}{p}a_{i}\in\Lambda$ for both $i=1,2$.
We set
$\mathcal{K}_{\neq 0}$ $=\left\{(a_{1},a_{2})\in\Lambda^{2}\colon 0\neq\det(a_{1})=\det(a_{2}),\,a_{1},a_{2}\text{ pair-primitive}\right\},$
$\mathcal{K}_{D}$ $=\{(a_{1},a_{2})\in\mathcal{K}_{\neq 0}\colon\det(a_{1})=-D\}.$
Note that $\operatorname{SL}_{2}(\mathbb{Z})\times\operatorname{SL}_{2}(\mathbb{Z})$ acts on these sets by individual conjugation. The aim of this subsection is to show that $\mathcal{K}_{D}$ parametrizes elements of $\mathcal{L}_{D}$.
Proposition 3.2 (The Klein map).
Consider the map
$\Phi:\mathcal{L}_{\neq 0}\to\Lambda\times\Lambda,\,L\mapsto(a_{1}(L),a_{2}(L)).$
The following holds:
  1. (1)
    For $L\in\mathcal{L}_{D}$ and for $i=1,2$ we have $\det(a_{i}(L))=-\operatorname{disc}([q_{L}])=-D$.
  2. (2)
    The map $\Phi$ is well-defined. Furthermore, we have
    $\Phi(L_{\mathrm{op}})=(-a_{1}(L),-a_{2}(L))=-\Phi(L).$
  3. (3)
    The map $\Phi$ induces bijections $\mathcal{L}_{\neq 0}\to\mathcal{K}_{\neq 0}$ and $\mathcal{L}_{D}\to\mathcal{K}_{D}$ for any discriminant $D\neq 0$. Moreover, the inverse is realized by $\Psi\colon\mathcal{K}_{\neq 0}\to\mathcal{L}_{\neq 0}$ given by
    (3.6) $(a_{1},a_{2})\mapsto L_{a_{1},a_{2}}:=\left\{x\in\operatorname{M}_{2}(\mathbb{Z}):a_{1}x=xa_{2}\right\}$
    where the orientation agrees with a basis of $L_{a_{1},a_{2}}\otimes\mathbb{Q}$ of the form $(a_{1}g,g)$ for some $0\neq g\in L_{a_{1},a_{2}}\otimes\mathbb{Q}$ with $\det(g)>0$, or with a basis of the form $(g,a_{1}g)$, for some $g\in L_{a_{1},a_{2}}\otimes\mathbb{Q}$ with $\det(g)<0$ (at least one of these options is realizable).
  4. (4)
    For $L\in\mathcal{L}_{\neq 0}$ with Klein vectors $a_{1},a_{2}$ we have $L^{\perp}=L_{-a_{1},a_{2}}$ or $L^{\perp}=L_{a_{1},-a_{2}}$ depending on the choice of orientation. In particular, $L^{\perp}\in\mathcal{L}_{D}$ whenever $L\in\mathcal{L}_{D}$.
  5. (5)
    The map $\Phi$ is $\operatorname{SL}_{2}(\mathbb{Z})\times\operatorname{SL}_{2}(\mathbb{Z})$-equivariant.
At this point we come back and define the orthogonal complement as an oriented plane as follows: for $(a_{1},a_{2})\in\mathcal{K}_{\neq 0}$
(3.7) $(L_{a_{1},a_{2}})^{\perp}=L_{-a_{1},a_{2}}.$
Remark 3.3.
The choice of orientation on the orthogonal complement in (3.7) is only one possible choice. Alternatively, one may choose an orientation on $L^{\perp}$ such that its oriented basis together with the oriented basis of $L$ forms an oriented basis of $\mathbb{R}^{4}$. These two choices only agree for planes $L$ with $\operatorname{disc}(q_{L})>0$. The choice here simplifies the exposition and is natural as it is induced by a linear automorphism on $\bigwedge^{2}\mathbb{Z}^{4}$ (compare this to the notion of the Hodge star operator).
Remark 3.4.
Formulas similar to (3.6) appear when exhibiting actions of the class group on e.g. integer points of the sphere – see for instance [11] and the references therein.
Proof of Proposition 3.2.
For item 1 we calculate:
$a_{1}(L)\overline{a_{1}(L)}$ $=\left(2v_{1}\overline{v_{2}}-\operatorname{Tr}\left(v_{1}\overline{v_{2}}\right)I_{2}\right)\left(2v_{2}\overline{v_{1}}-\operatorname{Tr}\left(v_{1}\overline{v_{2}}\right)I_{2}\right)$
$=\left(4\det(v_{1})\det(v_{2})I_{2}-2\operatorname{Tr}\left(v_{1}\overline{v_{2}}\right)\left(v_{1}\overline{v_{2}}+\overline{v_{1}\overline{v_{2}}}\right)+\left(\operatorname{Tr}\left(v_{1}\overline{v_{2}}\right)\right)^{2}I_{2}\right)$
$=\left(Q(v_{1})Q(v_{2})-\left(\operatorname{Tr}\left(v_{1}\overline{v_{2}}\right)\right)^{2}\right)I_{2}.$
Hence,
$\det(a_{1})(L)=\frac{1}{2}Q\left(a_{1}(L)\right)=\frac{1}{2}\operatorname{Tr}{a_{1}(L)\overline{a_{1}(L)}}=Q(v_{1})Q(v_{2})-\left(\operatorname{Tr}\left(v_{1}\overline{v_{2}}\right)\right)^{2}.$
The same calculation for $Q(a_{2}(L))$ yields the same expression, so $Q(a_{1}(L))=Q(a_{2}(L))$. Moreover, by writing out $q_{L}(x,y)$ as in (3.3) with the basis $(v_{1},v_{2})$ we see that
$-\operatorname{disc}q_{L}=4\det(v_{1})\det(v_{2})-Q(v_{1},v_{2})^{2}$ $=Q(v_{1})Q(v_{2})-\left(\operatorname{Tr}\left(v_{1}\overline{v_{2}}\right)\right)^{2}$
(3.8) $=\det(a_{i}(L))$
for $i=1,2$, as needed.
For item (2) we use (3.5) to extend $\Phi$ to any pair of linearly independent vectors $(v_{1},v_{2})\in\operatorname{M}_{2}(\mathbb{Q})^{2}$. In particular, for $L\in\mathcal{L}_{D}$, $\Phi(L)$ is $\Phi(v_{1},v_{2})$ for a basis $(v_{1},v_{2})$ of $L$. Note that $\Phi$ is multilinear and antisymmetric in $(v_{1},v_{2})$. It follows that calculating the Klein vectors using a basis of the form $(av_{1}+bv_{2},cv_{1}+dv_{2})$ multiplies both Klein vectors of $L$ by $\det\big(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\big)$, so item (2) follows. Moreover, in preparation for the proof of item 3 below, we observe that this also shows that if $\mathrm{span}(v_{1},v_{2})=\mathrm{span}(u_{1},u_{2})$, then $\lambda\Phi(v_{1},v_{2})=\Phi(u_{1},u_{2})$ for some $\lambda\in\mathbb{Q}^{\times}$ and the orientations of $(v_{1},v_{2})$ and $(u_{1},u_{2})$ agree if and only if $\lambda>0$.
We start the proof of item (3) by showing the containment $\Phi(\mathcal{L}_{D})\subset\mathcal{K}_{D}$. With item (1) pair-primitivity is all we need to prove. This can be proven similarly to [1, Lemma 2.4]; we give a variant thereof. As the map $\Phi$ is antisymmetric when extended to pairs of vectors as above, it factors through $\bigwedge^{2}\operatorname{M}_{2}$:
$\mathcal{L}$$\bigwedge^{2}\operatorname{M}_{2}(\mathbb{Z})$$\operatorname{M}_{2}^{0}(\mathbb{Z})\times\operatorname{M}_{2}^{0}(\mathbb{Z})$$\scriptstyle{\tilde{\Phi}}$
Denote by $(b_{1},b_{2},b_{3},b_{4})$ the basis $\mathcal{B}$ fixed in §3.1 and consider the basis $(c_{1}=b_{1}-b_{2},c_{2}=2b_{3},c_{3}=2b_{4})$ of $\Lambda$. Representing $\tilde{\Phi}$ with respect to the bases
$(b_{1}\wedge b_{2},b_{1}\wedge b_{3},b_{1}\wedge b_{4},b_{2}\wedge b_{3},b_{2}\wedge b_{4},b_{3}\wedge b_{4})\text{ and }$
$((c_{1},0),(c_{2},0),(c_{3},0),(0,c_{1}),(0,c_{2}),(0,c_{3}))$
we get the matrix
$\begin{pmatrix}1&0&0&0&0&-1\\ 0&0&0&-1&0&0\\ 0&0&-1&0&0&0\\ 1&0&0&0&0&1\\ 0&-1&0&0&0&0\\ 0&0&0&0&-1&0\end{pmatrix}$
which has determinant 2-2. This shows that $\tilde{\Phi}$ yields a bijection between primitive vectors in $\bigwedge^{2}\operatorname{M}_{2}(\mathbb{Z})$ and primitive vectors in the index two sublattice
$\Lambda^{\prime}=\{(a_{1},a_{2})\in\Lambda\times\Lambda:a_{1}+a_{2}\in 2\operatorname{M}_{2}(\mathbb{Z})\}$
of $\Lambda\times\Lambda$. It remains to show that the set $\mathcal{K}^{\prime}$ of primitive vectors $(a_{1},a_{2})$ in $\Lambda^{\prime}$ such that $\det(a_{1})=\det(a_{2})\neq 0$ is equal to $\mathcal{K}_{\neq 0}$. Clearly, if $(a_{1},a_{2})\in\mathcal{K}_{\neq 0}$ then the fact that $\det(a_{1})=\det(a_{2})$ implies that $a_{1}+a_{2}\in 2\operatorname{M}_{2}(\mathbb{Z})$ because $(a_{1})_{11},(a_{2})_{11}$ need to share the same parity. Also, $(a_{1},a_{2})$ is primitive in $\Lambda^{\prime}$ because it is primitive in the larger lattice $\Lambda\times\Lambda$. Thus, $\mathcal{K}_{\neq 0}\subset\mathcal{K}^{\prime}$. On the other hand, if $(a_{1},a_{2})\in\mathcal{K}_{\neq 0}$ were a non-trivial multiple of $(a_{1}^{\prime},a_{2}^{\prime})\in\Lambda^{2}$ then $2(a_{1}^{\prime},a_{2}^{\prime})=(a_{1},a_{2})$. But $a_{1}^{\prime},a_{2}^{\prime}$ have the same determinant and hence $a_{1}^{\prime}+a_{2}^{\prime}\in 2\operatorname{M}_{2}(\mathbb{Z})$ so that $(a_{1}^{\prime},a_{2}^{\prime})\in\Lambda^{\prime}$. This finishes the claims that $\mathcal{K^{\prime}}=\mathcal{K}_{\neq 0}$ and $\Phi(\mathcal{L}_{D})\subset\mathcal{K}_{D}$. We remark that the above claim is easier to prove on $\mathcal{L}_{D}$ when $D\equiv 1\,\mathrm{mod}\,4$.
For the reverse containment, we need first to discuss the inverse $\Psi$. Let $(a_{1},a_{2})\in\mathcal{K}_{\neq 0}$. Since $\det(a_{1})=\det(a_{2})$ and $a_{1},a_{2}$ are traceless, they have the same characteristic polynomial $x^{2}-D$. Thus, there exists $g\in\operatorname{M}_{2}^{\times}(\mathbb{Q})=\operatorname{GL}_{2}(\mathbb{Q})$ such that $ga_{2}g^{-1}=a_{1}$ (both $a_{1},a_{2}$ are conjugate to the companion matrix of $x^{2}-D$). In particular, $\Psi(a_{1},a_{2})=L_{a_{1},a_{2}}\neq\{0\}$. Since the subspace $L_{a_{1},a_{2}}(\mathbb{Q}):=L_{a_{1},a_{2}}\otimes\mathbb{Q}$ satisfies $L_{a_{1},a_{2}}(\mathbb{Q})=L_{a_{1},a_{1}}(\mathbb{Q})g$ and the centralizer $L_{a_{1},a_{1}}(\mathbb{Q})=\mathbb{Q}(a_{1})$ of $a_{1}$ has dimension $2$, $L_{a_{1},a_{2}}$ has rank $2$. It also follows that $(a_{1}g,g)$ is a basis of $L_{a_{1},a_{2}}(\mathbb{Q})$ as a $\mathbb{Q}$-vector space. We calculate $\Phi(a_{1}g,g)$:
(3.11) $\begin{array}[]{l}2a_{1}g\overline{g}-\operatorname{Tr}(a_{1}g\overline{g})I_{2}=2\det(g)\cdot a_{1}-0=2\det(g)\cdot a_{1}\\ 2\overline{g}a_{1}g-\operatorname{Tr}(\overline{g}a_{1}g)I_{2}=2\overline{g}ga_{2}-0=2\det(g)\cdot a_{2}.\end{array}$
By the claim from the end of the paragraph on item (2), we see that the choice of orientation in the statement of the proposition gives indeed a well-defined orientation on $L_{a_{1},a_{2}}$, so $\Psi(a_{1},a_{2})\in\mathcal{L}_{\neq 0}$ and $\Psi$ is well-defined.
A direct calculation shows that $v_{1},v_{2}\in L_{a_{1}(L),a_{2}(L)}$ so $L=L_{a_{1}(L),a_{2}(L)}$ as non-oriented subspaces. The observation in the proof of item (2) together with the calculation (3.11) upgrades this to an equality of planes. This concludes that $\Psi\circ\Phi|_{\mathcal{L}_{D}}=\operatorname{Id}_{\mathcal{L}_{D}}$. To see that $\Phi\circ\Psi|_{\mathcal{K}_{D}}=\operatorname{Id}_{\mathcal{K}_{D}}$, let $(a_{1},a_{2})\in\mathcal{K}_{D}$ and consider the following two bases for $L_{a_{1},a_{2}}$: a basis $(w_{1},w_{2})$ which spans $L_{a_{1},a_{2}}$ and a basis $(a_{1}g,g)$ of $L_{a_{1},a_{2}}(\mathbb{Q})$ as above (assume that there exist an element $g\in L_{a_{1},a_{2}}(\mathbb{Q})$ with $\det(g)>0$; the other case is dealt with similarly). We denote $(b_{1},b_{2}):=\Phi(L_{a_{1},a_{2}})=\Phi(w_{1},w_{2})$. By the observation above and (3.11), there exists $\lambda>0$ with
$(b_{1},b_{2})=\Phi(w_{1},w_{2})=\lambda\Phi(a_{1}g,g)=2\lambda\det(g)\cdot(a_{1},a_{2}).$
Since both $(b_{1},b_{2})$ and $(a_{1},a_{2})$ are primitive vectors in $\Lambda^{2}$ and $\lambda\det(g)>0$, they must be equal, which gives $\Phi\circ\Psi|_{\mathcal{K}_{D}}=\operatorname{Id}_{\mathcal{K}_{D}}$, completing the proof of item (3) .
For item (4) let $L\in\mathcal{L}_{D}$ and denote $a_{i}=a_{i}(L)$. Note that by item (3) we have that
(3.12) $L=\left\{x\in\operatorname{M}_{2}(\mathbb{Z}):a_{1}x=xa_{2}\right\}$
and that $\left\{y\in\operatorname{M}_{2}(\mathbb{Z}):-a_{1}y=ya_{2}\right\}$ is a plane in $\mathcal{L}_{D}$. We therefore show that, if $y\in\operatorname{M}_{2}(\mathbb{Q})$ with $a_{1}y=ya_{2}$ and $x\in\operatorname{M}_{2}(\mathbb{Q})$ with $a_{1}x=xa_{2}$, then $Q(x,y)=0$. Using that $a_{i}\overline{a_{i}}=\det(a_{i})\cdot I_{2}$ for $i=1,2$, we have
$\det(a_{1})\cdot Q(x,y)=\det(a_{1})\cdot\operatorname{Tr}(x\bar{y})=\operatorname{Tr}(\bar{a_{1}}a_{1}x\bar{y})=\operatorname{Tr}(a_{1}x\bar{y}\bar{a_{1}})=\\ =\operatorname{Tr}(a_{1}x\overline{\left(a_{1}y\right)})=\operatorname{Tr}(xa_{2}\overline{\left(-ya_{2}\right)})=\operatorname{Tr}(-xa_{2}\bar{a_{2}}\bar{y})=-\det(a_{2})\cdot Q(x,y).$
By item (1), $\det(a_{1})=\det(a_{2})\neq 0$, thus $Q(x,y)=0$ as desired.
Item (5) is a direct consequence of the definition of $a_{1}(\cdot)$ and $a_{2}(\cdot)$. ∎
3.3. Symplectic planes and their complements
We call a plane $L$ symplectic if we can find a basis $(v_{1},v_{2})$ of $L$ with $\theta(v_{1},v_{2})=1$. Such a basis will also be called symplectic. For a symplectic plane $L$ we define the symplectic complement $L^{\perp\kern-4.2pt\perp}$ as
$L^{\perp\kern-4.2pt\perp}=\left\{x\in\operatorname{M}_{2}(\mathbb{Z})\colon\theta(x,\ell)=0,\,\forall\ell\in L\right\}.$
where the orientation is given by any symplectic basis.
The following observation (3.13) is elementary, but it will be essential for relating information about $L^{\perp}$ to information about the symplectic complement $L^{\perp\kern-4.2pt\perp}$. Using the explicit forms of $\theta$ and $Q$ as in (3.2) one can check directly that, a priori only as non-oriented planes, we have
(3.13) $L^{\perp\kern-4.2pt\perp}=\left(Lj\right)^{\perp}=L^{\perp}\overline{j}=L^{\perp}(-j)=L^{\perp}j$
where we view $L,L^{\perp},L^{\perp\kern-4.2pt\perp}$ as subsets of $\operatorname{M}_{2}(\mathbb{Z})$. We further have for $v,w\in\operatorname{M}_{2}(\mathbb{Z})$
(3.14) $Q(vj,wj)=\operatorname{Tr}\left(vj\overline{j}\overline{w}\right)=-\operatorname{Tr}\left(v\overline{w}\right)=-Q(v,w),$
and a similar calculation gives
(3.15) $\theta(vj,wj)=\operatorname{Tr}\left(vjj\overline{j}\overline{w}\right)=-\operatorname{Tr}\left(vj\overline{w}\right)=-\theta(v,w).$
Lemma 3.5.
A plane $L\in\mathcal{L}$ is symplectic if and only if $a_{2}(L)=\big(\begin{smallmatrix}1&\alpha\\ \gamma&-1\end{smallmatrix}\big)$ for some $\alpha,\gamma\in 2\mathbb{Z}$.
Proof.
First note that we already know from Proposition 3.2 that $a_{2}(L)$ has the form $\big(\begin{smallmatrix}\beta&\alpha\\ \gamma&-\beta\end{smallmatrix}\big)$ for $\beta\in\mathbb{Z},\alpha,\gamma\in 2\mathbb{Z}$. By (3.2) we have
$\theta(v_{1},v_{2})=\operatorname{Tr}\left(v_{1}j\overline{v_{2}}\right)=\operatorname{Tr}\left(\overline{v_{2}}v_{1}j\right)=\operatorname{Tr}\big(\tfrac{1}{2}a_{2}(L)j\big)=\beta$
where the second to last equality follows as $\operatorname{Tr}\left(I_{2}j\right)=0$. The lemma follows. ∎
Lemma 3.5 together with the previous observations enable us to relate the Klein vectors and the associated quadratic forms for $L,\,L^{\perp}$ and $L^{\perp\kern-4.2pt\perp}$:
Proposition 3.6.
Let $L\in\mathcal{L}_{D}$ be a symplectic plane. Then $L^{\perp}$ is symplectic. Furthermore, $L^{\perp\kern-4.2pt\perp}\in\mathcal{L}_{D}$ and $q_{L^{\perp\kern-4.2pt\perp}}]=[-\overline{q_{L^{\perp}}}]$. If we write $a_{2}(L)=\big(\begin{smallmatrix}1&\alpha\\ \gamma&-1\end{smallmatrix}\big)$, then
(3.16) $\Phi(L^{\perp\kern-4.2pt\perp})=(-a_{1}(L),\big(\begin{smallmatrix}1&-\alpha\\ -\gamma&-1\end{smallmatrix}\big)).$
Proof.
Recall that $L^{\perp}=\Psi(-a_{1}(L),a_{2}(L))$ and so $L^{\perp}$ is symplectic by Lemma 3.5. Choose then a symplectic basis $(w_{1},w_{2})$ for $L^{\perp}$ and notice that $(w_{2}j,w_{1}j)$ is a (a priori non-oriented) basis of $L^{\perp\kern-4.2pt\perp}$ by (3.13). By (3.15), $\theta(w_{2}j,w_{1}j)=-\theta(w_{2},w_{1})=1$ so it is a symplectic basis. We can therefore use it to calculate the quadratic form on $L^{\perp\kern-4.2pt\perp}$ as well as $\Phi(L^{\perp\kern-4.2pt\perp})$. For the former, note that
$[q_{L^{\perp\kern-4.2pt\perp}}]=[\det(xw_{2}j+yw_{1}j)]=[-\det(xw_{2}+yw_{1})]=[-\overline{q_{L^{\perp}}}].$
In particular, the statement about the discriminant of $L^{\perp\kern-4.2pt\perp}$ follows. It remains to compute the Klein vectors:
(3.17) $a_{1}(L^{\perp\kern-4.2pt\perp})=2w_{2}j\overline{w_{1}j}-\operatorname{Tr}\left(w_{2}j\overline{w_{1}j}\right)I_{2}=-2w_{2}\overline{w_{1}}+\operatorname{Tr}\left(w_{2}\overline{w_{1}}\right)I_{2}=\\ =-\overline{a_{1}(L^{\perp})}=a_{1}(L^{\perp})=-a_{1}(L)$
and
(3.18) $a_{2}(L^{\perp\kern-4.2pt\perp})=2\overline{w_{1}j}w_{2}j-\operatorname{Tr}\left(\overline{w_{1}j}w_{2}j\right)I_{2}=-2j\overline{w_{1}}w_{2}j+\operatorname{Tr}\left(j\overline{w_{1}}w_{2}j\right)I_{2}=\\ =-j\overline{a_{2}(L^{\perp})}j=j{a_{2}(L^{\perp})}j=\big(\begin{smallmatrix}1&-\alpha\\ -\gamma&-1\end{smallmatrix}\big),$
where we used in the last equality that $a_{2}(L^{\perp})=a_{2}(L)$. ∎
4. Gauss composition via Klein vectors and the Bhargava cube
In this section, we establish the expression for Gauss composition in terms of Klein vectors as announced in the introduction in Theorem 1.6. For readability, we first do the case where all binary quadratic forms are primitive. We then prove the more general case where the condition of primitivity is weakened to pair-primitivity. The argument in this general case is similar to that of the primitive case, but requires more care, as classically, Gauss composition is only defined for primitive forms and for the non-primitive case, more setup is needed.
To understand how the Klein vectors $a_{1}(L),a_{2}(L)$ determine the quadratic form $q_{L}]$, we identify them with IBQFs via the map $A\in\Lambda\mapsto q_{A}$ from §2. Two vectors in the Gross lattice $\Lambda$ are pair-primitive (cf. Definition 3.1) if and only if the associated IBQFs have coprime content.
4.1. The primitive case
Recall that $\mathcal{L}_{\neq 0}$ denotes the set of planes $L$ with $\operatorname{disc}(q_{L})$ non-zero. The following is the primitive version of Theorem 4.5 below.
Theorem 4.1.
Let $L=L_{a_{1},a_{2}}\in\mathcal{L}_{\neq 0}$ and assume that $a_{1},a_{2}\in\Lambda$ are both primitive. Then $q_{L}$ is primitive, and we have
(4.1) $q_{L}]=[q_{a_{1}}]^{-1}*[q_{a_{2}}].$
Proof using Dirichlet composition.
We refer to e.g. [7, Ch. 14] for general facts regarding Dirichlet composition. Using equivariance of the Klein map $\Phi$ (see Proposition 3.2(5)) as well as the map $A\mapsto q_{A}$ (see §2) we may bring $\overline{q_{a_{1}}},q_{a_{2}}$ into concordant form (cf. [7, p. 335]) without changing $q_{L}]$. That is, we suppose that $\overline{q_{a_{1}}}=\alpha_{1}x^{2}+\beta xy+\gamma_{1}y^{2}$ and $q_{a_{2}}=\alpha_{2}x^{2}+\beta xy+\gamma_{2}y^{2}$ where $\alpha_{1},\alpha_{2}$ are coprime. Since the discriminants of $q_{a_{1}},q_{a_{2}}$ agree, we have $\alpha_{1}\gamma_{1}=\alpha_{2}\gamma_{2}$ and thus $\alpha_{1}\mid\gamma_{2}$ and $\alpha_{2}\mid\gamma_{1}$. By the explicit inverse in Proposition 3.2 we have
$L=\left\{x\in\operatorname{M}_{2}(\mathbb{Z}):\big(\begin{smallmatrix}-\beta&-2\alpha_{1}\\ 2\gamma_{1}&\beta\end{smallmatrix}\big)x=x\big(\begin{smallmatrix}\beta&-2\alpha_{2}\\ 2\gamma_{2}&-\beta\end{smallmatrix}\big)\right\}.$
It is readily checked that
$v_{1}=\big(\begin{smallmatrix}-\alpha_{1}&0\\ \beta&-\alpha_{2}\end{smallmatrix}\big),\ v_{2}=\big(\begin{smallmatrix}0&-1\\ \frac{\gamma_{2}}{\alpha_{1}}&0\end{smallmatrix}\big)$
is an oriented basis of $L$. Since $\det(xv_{1}+yv_{2})=\alpha_{1}\alpha_{2}x^{2}+\beta xy+\star y^{2}$, the definition of the Gauss composition via concordant forms implies that $q_{L}]$ is indeed the composition of $q_{a_{1}}]^{-1}$ and $q_{a_{2}}]$. ∎
We provide a second proof using Bhargava’s cube law. This proof has the advantage that we do not need to choose special representatives of the classes of the involved quadratic forms: it works directly with the given plane $L$ rather than the equivalence class of $L$ under $\operatorname{SL}_{2}(\mathbb{Z})\times\operatorname{SL}_{2}(\mathbb{Z})$. As a by-product, this proof provides a recipe to construct a Bhargava cube that realizes two given quadratic forms as slicings of the cube; see Remark 4.10.
Proof using the Bhargava cube.
Take a basis $(v_{1},v_{2})$ of $L(\mathbb{Z})$ with $\Phi(v_{1},v_{2})=(a_{1},a_{2})$ and set $M_{1}=v_{1},\,N_{1}=v_{2}$ in the notation of [3, §2.1] to define a Bhargava cube. The cube law yields three IBQFs $q_{1},q_{2},q_{3}$ with $q_{1}]*[q_{2}]*[q_{3}]=\operatorname{Id}$. By definition of $M_{1}$ and $N_{1}$, $q_{1}=-\det(xM_{1}-yN_{1})=-q_{L}(-x,y)=-\overline{q_{L}}(x,y)$, and a direct calculation shows that $q_{2}=\big(\begin{smallmatrix}0&1\\ 1&0\end{smallmatrix}\big).q_{a_{2}}$ and $q_{3}=-\overline{q_{a_{1}}}$. As $q_{a_{1}}$ and $q_{a_{2}}$ are assumed to be primitive, $q_{2}$ and $q_{3}$ are primitive. It follows then that $q_{1}$ and therefore $q_{L}$ is primitive (use for example the equivalence of the cube laws with Dirichlet’s composition, see [3, Equation (41)]). Moreover, we have $\overline{q_{L}}]*[\overline{q_{a_{2}}}]*[-\overline{q_{a_{1}}}]=\mathrm{id}$ or equivalently (as $\overline{\cdot}$ is inversion and the class group is abelian)
(4.2) $[-\overline{q_{L}}]=[q_{a_{2}}]*[-q_{a_{1}}]=[-q_{a_{1}}]*[q_{a_{2}}].$
It remains to show that this implies (4.1). To that end, we only need to prove that for any two primitive IBQFs $q_{1}$ and $q_{2}$ of the same discriminant we have $\overline{[q_{1}]*[q_{2}]}=[-\overline{q_{1}}]*[q_{2}]$. This can be either done using the Bhargava cube (cf. Lemma 4.7) or as follows: assume that $q_{1}$ and $q_{2}$ are concordant, i.e. $q_{1}=a_{1}x^{2}+bxy+c_{1}y^{2}$ and $q_{2}=a_{2}x^{2}+bxy+c_{2}y^{2}$ with $a_{1}a_{2}\neq 0$ and such that $a_{1}$ and $a_{2}$ are coprime (see [7, p. 336]). Then their composition is $q_{3}(x,y)=a_{1}a_{2}x^{2}+bxy+\frac{D-b^{2}}{4a_{1}a_{2}}y^{2}$ (see [7, p. 335]). Note that $\overline{q_{1}}$ and $q_{2}$ are also concordant, and their composition is $\overline{q_{3}}$. This proves the above claim and hence the theorem. ∎
4.2. The general case
Previously, we have only discussed Gauss composition for primitive quadratic forms of equal discriminant.
In general, there is no unique composition of classes of non-primitive quadratic forms. However, in the case where two classes $q_{1}]$, $q_{2}]$ have coprime content, it is known that the composition $q_{1}][q_{2}]$ coming from the Dirichlet law is well-defined. This was written down first in [24], and generalized to quadratic forms over integral domains in [6]. For a more modern perspective, see Penner [25].
Proposition 4.2 ([24], Theorem 2).
Let $s_{1}$, $s_{2}$ be ideal classes of discriminant $D$, with content $m_{1}$, resp. $m_{2}$, such that $\gcd(m_{1},m_{2})=1$.
Then one can (non-uniquely) find a pair of concordant forms $q_{1}=a_{1}x^{2}+bxy+a_{2}cy^{2}\in s_{1}$ and $q_{2}=a_{2}x^{2}+bxy+a_{1}cy^{2}\in s_{2}$, and the class of
$q_{1}q_{2}=a_{1}a_{2}x^{2}+bxy+cy^{2}$
is independent of the choices of $q_{1}$ and $q_{2}$ and has content $m_{1}m_{2}$. That is, there is a well-defined composition law (known as Dirichlet composition)
$\{s_{1}\in\mathcal{Q}_{D}^{+}$ $:\mathrm{content}(s_{1})=m_{1}\}\times\{s_{2}\in\mathcal{Q}_{D}^{+}:\mathrm{content}(s_{2})=m_{2}\}$
$\longrightarrow\{s\in\mathcal{Q}_{D}^{+}:\mathrm{content}(s)=m_{1}m_{2}\},$
whenever $\gcd(m_{1},m_{2})=1$.
Proposition 4.3.
The above composition law is associative when defined.
Proof.
This is proved in [6, Thm. 3.5] in the general case where $\mathbb{Z}$ is replaced with a general integral domain. ∎
Corollary 4.4.
For any integer $m$ with $m^{2}\mid D$, the group $\mathcal{G}_{D}^{+}$ of equivalence classes of primitive IBQFs of discriminant $D$ acts on the set of equivalence classes of IBQFs of discriminant $D$ and content $m$ by composition. Moreover, there is a surjective homomorphism $\mathcal{G}_{D}^{+}\to\mathcal{G}_{D/m^{2}}^{+},t\mapsto\tilde{t}$ with the following compatibility rule: for any $t\in\mathcal{G}_{D}^{+}$ and for any $s\in\mathcal{Q}_{D}^{+}$ of content $m$ we have
(4.3) $\tfrac{1}{m}(t*s)=\tilde{t}*(\tfrac{1}{m}s).$
Proof.
The first part is immediate from Proposition 4.3. For the second part, the map associating to $t\in\mathcal{G}_{D}^{+}$ the class $\tilde{t}=\frac{1}{m}(t\ast(m\,1_{D/m^{2}}))\in\mathcal{G}_{D/m^{2}}^{+}$ satisfies the requirements – see [25, p. 2706]. ∎
We now have everything we need to state and prove the version of Theorem 4.1 without the primitivity assumption. This is equivalent to Theorem 1.6.
Theorem 4.5.
For any $L=L_{a_{1},a_{2}}\in\mathcal{L}_{\neq 0}$ with $a_{1},a_{2}\in\Lambda$ pair-primitive we have
$[q_{L}]=\overline{[q_{a_{1}}]}*[q_{a_{2}}].$
Moreover, the content of $q_{L}]$ is the product of the contents of $q_{a_{1}},q_{a_{2}}$. If $a_{1},a_{2}$ are primitive, then $q_{L}]=[q_{a_{1}}]^{-1}*[q_{a_{2}}]$ is primitive. In particular, this recovers the primitive case, i.e., Theorem 4.1.
Proof.
This is analogous to the proof of Theorem 4.1 except that one can no longer take the inverse of an ideal class. The proof using Dirichlet composition goes trough exactly as before. The proof using the Bhargava cube works the same after some small preparations, which we provide in Proposition 4.6 and Lemma 4.7. ∎
Proposition 4.6.
Let $A$ be a Bhargava cube producing a triple $(q_{1},q_{2},q_{3})$ of quadratic forms. Suppose that $\gcd(\operatorname{content}(q_{2}),\operatorname{content}(q_{3}))=1$. Then $q_{2}]*[q_{3}]=\overline{[q_{1}]}$.
Conversely, if $q_{2}]*[q_{3}]=\overline{[q_{1}]}$ and $\gcd(\operatorname{content}(q_{2}),\operatorname{content}(q_{3}))=1$ then there exists a Bhargava cube producing the triple of forms $(q_{1},q_{2},q_{3})$.
Proof.
This follows by the argument in the Appendix of [3] showing the equivalence of Bhargava’s cube law with Dirichlet composition. ∎
We now state a lemma about how this composition law interacts with the maps $q]\mapsto[\overline{q}]$ and $q]\mapsto[-q]$.
Lemma 4.7.
If $q_{1}]=[q_{2}]*[q_{3}]$, then
  1. a)
    $\overline{q_{1}}]=[\overline{q_{2}}]*[\overline{q_{3}}]$
  2. b)
    $q_{1}]=[\overline{q_{2}}]*[-q_{3}]=[-q_{2}]*[\overline{q_{3}}]$
  3. c)
    $\overline{q_{1}}]=[q_{2}]*[-\overline{q_{3}}]=[-\overline{q_{2}}]*[q_{3}]$
Note that a version of the lemma was already used and proven in the proof of Theorem 4.1. Here, we give a different proof using the Bhargava cube.
Proof.
Part a) follows from the fact that reflecting a Bhargava cube around its center produces another Bhargava cube.
For part b), observe that multiplying one of two layers of a Bhargava cube by 1-1 replaces the quadratic form $q_{i}$ coming from that splitting by $\overline{q_{i}}$, while replacing the quadratic forms coming from the other two splittings by their negatives.
Doing this to the Bhargava cube producing the triple $(\overline{q_{1}},q_{2},q_{3})$, we get a Bhargava cube producing $(-\overline{q_{1}},\overline{q_{2}},-q_{3})$, so $q_{1}]=[\overline{q_{2}}]*[-q_{3}]$, and the other result follows symmetrically.
Finally, part c) follows from applying part a) to the equalities in part b). ∎
4.3. Consequences for $L^{\perp}$ and $L^{\perp\kern-4.2pt\perp}$
We now deduce expressions for $q_{L^{\perp}}$ and $q_{L^{\perp\kern-4.2pt\perp}}$ using Theorem 4.5.
Corollary 4.8.
For any $L=L_{a_{1},a_{2}}\in\mathcal{L}_{\neq 0}$ we have $q_{L^{\perp}}]=\overline{[-q_{a_{1}}]}*[q_{a_{2}}]$. Furthermore, if $L$ is symplectic,
$[q_{L^{\perp\kern-4.2pt\perp}}]$ $=[-\overline{q_{L^{\perp}}}]=[q_{a_{1}}]*[q_{a_{2}}].$
Moreover, for primitive $a\in\Lambda$, we have $q_{\langle\mathrm{id},a\rangle^{\perp}}]=-[q_{a}]^{2}$.
In Corollary 4.8, $\langle\mathrm{id},a\rangle^{\perp}$ is taken to be (the integer points of) the orthogonal complement of the plane $\langle\mathrm{id},a\rangle\cap\operatorname{M}_{2}(\mathbb{Z})$ with orientation given by $\mathrm{id},a$.
Proof.
The formula for $q_{L^{\perp}}]$ follows from applying Theorem 4.5 to $L^{\perp}{=}L_{-a_{1},a_{2}}$ (see (3.7)) and using $q_{-a_{1}}=-q_{a_{1}}$. The formulas for $q_{L^{\perp\kern-4.2pt\perp}}]$ are obtained from Proposition 3.6 and Lemma 4.7. Alternatively, one may apply the formulas in Proposition 3.6 for $a_{1}(L^{\perp\kern-4.2pt\perp})$ and $a_{2}(L^{\perp\kern-4.2pt\perp})$ in conjunction with Theorem 4.5. The last equality follows from $\langle\mathrm{id},a\rangle^{\perp}=(L_{-a,-a})^{\perp}=L_{a,-a}$ (cf. (3.7)) and Lemma 4.7b). ∎
For a symplectic subspace $L=L_{a_{1},a_{2}}$ we have by Lemma 3.5 that $a_{2}$ is automatically primitive and $q_{a_{2}}]\in\mathcal{G}_{D}^{+}$. We therefore have the following.
Corollary 4.9.
For a symplectic subspace $L=L_{a_{1},a_{2}}$ we have
$q_{L}]=[q_{L^{\perp\kern-4.2pt\perp}}]\iff[q_{a_{1}}]=\overline{[q_{a_{1}}]},$
and
$q_{L}]=\overline{[q_{L^{\perp\kern-4.2pt\perp}}]}\iff[q_{a_{2}}]^{2}\in\operatorname{Stab}_{\mathcal{G}_{D}^{+}}([q_{L}])\iff-[q_{a_{2}^{\perp}}]\in\operatorname{Stab}_{\mathcal{G}_{D}^{+}}([q_{L}]).$
where $\operatorname{Stab}_{\mathcal{G}_{D}^{+}}([q_{L}])$ is the set of elements $g$ in the group $\mathcal{G}_{D}^{+}$} such that $g[q_{L}]=[q_{L}]$. In particular, if $q_{L}$ is primitive then
$[q_{L}]\neq[q_{L^{\perp\kern-4.2pt\perp}}]\iff[q_{a_{1}}]^{2}\neq 1,\quad[q_{L}]\neq\overline{[q_{L^{\perp\kern-4.2pt\perp}}]}\iff[q_{a_{2}}]^{2}\neq 1.$
Remark 4.10.
Given two IBQF’s $q_{1},q_{2}$ the proofs of Theorems 4.1 and 4.5 show how one can find a Bhargava cube such that $q_{1},q_{2}$ arise from slicings of that cube. Take $a_{1}=A(q_{1})$, $a_{2}=-A(\big(\begin{smallmatrix}0&-1\\ 1&0\end{smallmatrix}\big).q_{2})$ and compute the plane $L=L_{a_{1},a_{2}}$. Given any basis $v_{1},v_{2}$ of $L$, the cube with $M_{1}=v_{1},\,N_{1}=v_{2}$ (cf. [3, §2.1]) is such that the given forms $q_{1},q_{2}$ are obtained from slicings as desired.
Remark 4.11 (Relations of IBQFs).
In the above (Theorem 4.5 and Corollary 4.8), we have established exact relations in the six tuple
$\big([q_{a_{1}}],[q_{a_{2}}],$ $[q_{L}],[q_{L^{\perp}}],[q_{\langle\mathrm{id},a_{1}\rangle^{\perp}}],[q_{\langle\mathrm{id},a_{2}\rangle^{\perp}}]\big)$
(4.4) $=\big([q_{a_{1}}],[q_{a_{2}}],\overline{[q_{a_{1}}]}*[q_{a_{2}}],\overline{[-q_{a_{1}}]}*[q_{a_{2}}],-[q_{a_{1}}]^{2},-[q_{a_{2}}]^{2}\big)$
for any $L=L_{a_{1},a_{2}}\in\mathcal{L}_{D}$. These are the tuples considered in the equidistribution results of [1].
5. Knot theory: realization of disjoint Seifert surfaces with prescribed Seifert forms
In this section we set up the relevant notions from knot theory and prove Theorems 1.1 and 1.2. In §5.1 we recall the notion of Seifert forms and check that for closed surfaces $\Sigma\subset S^{3}$ of genus $2$ the Seifert form on $H_{1}(\Sigma;\mathbb{Z})\cong\mathbb{Z}^{4}$ is, up to a choice of coordinates, the bilinear form $(\mathbb{Z}^{4},U)$ studied in §3. This is the reason that the form $U$ has relevance to the Seifert surface questions outlined in the beginning of the introduction. In §5.2 we show that $2$-dimensional symplectic planes $L$ of $H_{1}(\Sigma;\mathbb{Z})$ (with respect to the intersection form) correspond to the first homology of a genus $1$ surface $F\subset\Sigma$ whose boundary is a knot. Moreover, the symplectic complement $L^{\perp\kern-4.2pt\perp}$ of $L$ is equal to the first homology of the complement $\Sigma\setminus F$. We then show how the results about symplectic subspaces in $(\mathbb{Z}^{4},U)$ and their complements from §3 imply the Seifert form realization result as stated in Theorem 1.1. Finally, in §5.3, we describe the intersection form on second homology of the double branched cover of Seifert surfaces that are pushed into $B^{4}$. We then use the $3$-dimensional Seifert form realization results to establish the existence of many pairs of genus 1 Seifert surfaces that are not ambiently isotopic after being pushed into $B^{4}$. In particular, we prove Theorem 1.2.
5.1. Setup: Seifert form and Seifert surfaces
We point to [20, Chap. 2] for details on the Seifert form. For an oriented compact smooth surface with (possibly empty) boundary $F\subset S^{3}$, we denote by
$U_{F}\colon H_{1}(F;\mathbb{Z})\times H_{1}(F;\mathbb{Z})\to\mathbb{Z}$
the Seifert form. Representing elements in $H_{1}(F;\mathbb{Z})$ as curves, the Seifert form may be defined by $U_{F}([\alpha],[\beta])=\mathrm{lk}(\alpha,\beta^{+})$. Here, $\mathrm{lk}$ denotes Gauss’s linking number of disjoint curves in $S^{3}$, $\alpha$ and $\beta$ are curves in $F$ representing a given homology class, and $\beta^{+}$ is the result of pushing $\beta$ a small amount along a positive normal into $S^{3}\setminus F$; see the right-hand side of Figure 1.
$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\beta_{1}$}
$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\beta_{2}$}
$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\alpha_{1}$}
$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\alpha_{2}$}
$\alpha$
$\beta$
$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\beta^{+}$}
Figure 1. Left: the surface $\Sigma_{\rm std}$ (gray) with simple closed curves $\alpha_{1},\beta_{1},\alpha_{2},\beta_{2}$ (red) such that the corresponding homology classes form a basis. Right: Illustration of the positive push-off $\beta^{+}$ (red) of $\beta$ in $\Sigma$ (gray) needed for the definition of the Seifert form.
Writing $\overline{F}$ for the surface $F$ with reversed orientation, we observe that $U_{\overline{F}}(a,b)=U_{F}(b,a)$. Recall that the antisymmetrization of $U_{F}$,
$\theta_{F}\colon H_{1}(F;\mathbb{Z})\times H_{1}(F;\mathbb{Z})\to\mathbb{Z},\,(a,b)\mapsto U_{F}(a,b)-U_{F}(b,a)=U_{F}(a,b)-U_{\overline{F}}(a,b),$
is the intersection form on $H_{1}(F;\mathbb{Z})$. We denote by $Q_{F}$ the symmetrization of the Seifert form, i.e. $Q_{F}\colon H_{1}(F;\mathbb{Z})\times H_{1}(F;\mathbb{Z})\to\mathbb{Z},(a,b)\mapsto U_{F}(a,b)+U_{F}(b,a)$. We choose the notation $U_{F}$ for the Seifert form due to the following.
Example 5.1.
Let $\Sigma_{\rm std}$ be the unknotted, oriented genus two closed subsurface of $S^{3}$ as depicted in Figure 1. Then there exists a basis $\mathcal{B}=\left([\alpha_{1}],[\beta_{1}],[\alpha_{2}],[\beta_{2}]\right)$ of $H_{1}(\Sigma_{\rm std};\mathbb{Z})$, where $\alpha_{i}$ and $\beta_{i}$ are as indicated in Figure 1, such that with respect to this basis the Seifert form is represented by the matrix $U=\big(\begin{smallmatrix}0&1\\ 0&0\end{smallmatrix}\big)\oplus\big(\begin{smallmatrix}0&1\\ 0&0\end{smallmatrix}\big)$ from (3.1).
In fact, every closed oriented surface of genus $g$ has as its Seifert form the $g$-fold direct sum $\big(\begin{smallmatrix}0&1\\ 0&0\end{smallmatrix}\big)\oplus\cdots\oplus\big(\begin{smallmatrix}0&1\\ 0&0\end{smallmatrix}\big)$.
Lemma 5.2.
Let $\Sigma$ be an oriented closed genus $g$ surface in $S^{3}$. Then there is a basis of $H_{1}(\Sigma;\mathbb{Z})$ in which $U_{\Sigma}$ is represented by
$\underbrace{\big(\begin{smallmatrix}0&1\\ 0&0\end{smallmatrix}\big)\oplus\cdots\oplus\big(\begin{smallmatrix}0&1\\ 0&0\end{smallmatrix}\big)}_{g}.$
Proof.
Write $S^{3}=X\cup Y$, where $X$ and $Y$ are 3-manifolds such that $X\cap Y=\Sigma$ (as sets) and such that $X$ and $Y$ have oriented boundary $\overline{\Sigma}$ and $\Sigma$}, respectively. In particular, if $\gamma$ is a curve in $\Sigma$, then $\gamma^{+}$ is in the interior of $X$. Applying Mayer-Vietoris to this decomposition of $S^{3}$ implies that $H_{1}(\Sigma;\mathbb{Z})\cong H_{1}(X;\mathbb{Z})\oplus H_{1}(Y;\mathbb{Z})$, where the isomorphism is induced by the sum of the homomorphisms induced by inclusion, since $H_{2}(S^{3};\mathbb{Z})$ and $H_{1}(S^{3};\mathbb{Z})$ (these are the terms before $H_{1}(\Sigma;\mathbb{Z})$ and after $H_{1}(X;\mathbb{Z})\oplus H_{1}(Y;\mathbb{Z})$ in the long exact sequence) are trivial. Noting that $H_{1}(X;\mathbb{Z})$ and $H_{1}(Y;\mathbb{Z})$ have rank $g$ (e.g. by the “half lives, half dies” principle that says that the homomorphisms $H_{1}(\Sigma,\mathbb{Q})\to H_{1}(X,\mathbb{Q})$ and $H_{1}(\overline{\Sigma},\mathbb{Q})\to H_{1}(Y,\mathbb{Q})$ induced by the inclusions have rank $g$), there exists a basis
$([\alpha_{1}],[\alpha_{2}],\dots,[\alpha_{g}],[\beta_{1}],\dots,[\beta_{g}])$
of $H_{1}(\Sigma;\mathbb{Z})$ such that the $\alpha_{i}$ are closed curves that are null-homologous in $X$ and the $\beta_{i}$ are null-homologous in $Y$. Since primitive elements in $H_{1}(\Sigma;\mathbb{Z})$ are representable by simple closed curves we may and do assume that the $\alpha_{i}$ and $\beta_{i}$ are simple closed curves. We note that for a simple closed curve $\alpha$ that is null-homologous (i.e. the boundary of an oriented surface) in $X$ and a curve $\beta$ that is null-homologous in $Y$, one has $\mathrm{lk}(\alpha,\alpha^{+})=\mathrm{lk}(\beta,\beta^{+})=\mathrm{lk}(\beta,\alpha^{+})=0$. (For this recall that $\mathrm{lk}(\delta,\gamma)$ equals the algebraic intersection of $\gamma$ with an oriented surface with boundary $\delta$, and note that we can choose surfaces for $\alpha$, $\beta$, and $\beta$ that are disjoint from $\alpha^{+}$, $\beta^{+}$, and $\alpha^{+}$, respectively.) This implies that the matrix of $U_{\Sigma}$ with respect to this basis has the last $g$ rows and the first $g$ columns all zero. In other words, with respect to this basis, $U_{\Sigma}$ is given by $\big(\begin{smallmatrix}0&A\\ 0&0\end{smallmatrix}\big)$, where 0 denotes the $g\times g$-matrix with entries 0 and $A$ is a $g\times g$-matrix. Noting that $A$ is invertible (since the antisymmetrization $\big(\begin{smallmatrix}0&A\\ -A^{T}&0\end{smallmatrix}\big)$ represents the intersection form on $\Sigma$, which has determinant 1), we can and do assume (by applying a base change if needed) that $A$ is the identity matrix. This establishes that, with respect to the basis
$([\alpha_{1}],[\beta_{1}],[\alpha_{2}],[\beta_{2}],\dots,[\alpha_{g}],[\beta_{g}]),$
$U_{\Sigma}$ is, as desired, given by the matrix
$\underbrace{\big(\begin{smallmatrix}0&1\\ 0&0\end{smallmatrix}\big)\oplus\cdots\oplus\big(\begin{smallmatrix}0&1\\ 0&0\end{smallmatrix}\big)}_{g}.\qed$
Given Lemma 5.2, it is no surprise that most commonly the Seifert form is considered for surfaces with non-empty boundary. A Seifert surface $F$ (for a knot or link $K$) is an oriented compact non-empty smooth surface in $S^{3}$ without closed components (with oriented boundary $K$). If $F$ is a genus $1$ Seifert surface for a knot $K$, we write
(5.1) $s_{F}=[U_{F}(a,a)x^{2}+(U_{F}(a,b)+U_{F}(b,a))xy+U_{F}(b,b)y^{2}]\in\mathcal{Q}^{+}_{D}$
where $(a,b)$ is a symplectic basis for $H_{1}(F;\mathbb{Z})$ (i.e. a basis with respect to which the intersection form is given by $\big(\begin{smallmatrix}0&1\\ -1&0\end{smallmatrix}\big)$) and $D$ is the discriminant $(U_{F}(a,b)+U_{F}(b,a))^{2}-4U_{F}(a,a)U_{F}(b,b)$. Note that $s_{\overline{F}}=\overline{s_{F}}$ since $(a,b)$ is a symplectic basis for $H_{1}(F;\mathbb{Z})$ if and only if (a,b)(a,-b) is a symplectic basis for $H_{1}(\overline{F};\mathbb{Z})$.
5.2. Realization of Seifert forms and proof of Theorem 1.1
Lemma 5.3.
Let $\Sigma$ be an oriented closed smooth genus 2 surface in $S^{3}$. For every symplectic plane $L$ of $H_{1}(\Sigma;\mathbb{Z})$, there exists a non-contractible separating simple closed curve $K\subset\Sigma$ such that $L=H_{1}(F_{K};\mathbb{Z})\subset H_{1}(\Sigma;\mathbb{Z})$, where $F_{K}$ is the closure of the component of $\Sigma\setminus K$ that induces the correct orientation on $K$.
Proof.
Let $(a,b)$ be a symplectic basis of $L$; that is, $a$ and $b$ are primitive and satisfy $\theta(a,b)=1$. There exist simple closed curves $\alpha$ and $\beta$ that intersect once transversely such that $a=[\alpha]$ and $b=[\beta]$ (e.g. by [12, Theorem 6.4]). Consider a closed deformation retraction neighborhood $N$ of $\alpha\cup\beta$ that is a smooth surface with boundary. $N$ is diffeomorphic to a genus 1 surface with one boundary component. Set $K\coloneqq\partial N$ and $F_{K}\coloneqq N$. By construction $([\alpha],[\beta])$ forms a basis of $H_{1}(F_{K};\mathbb{Z})$, hence $L=H_{1}(F_{K};\mathbb{Z})$. ∎
Lemma 5.4.
Let $K$ be a non-contractible separating simple closed curve in a closed oriented smooth genus 2 surface $\Sigma$ such that $F_{K}$ is the closure of the component of $\Sigma\setminus K$ that induces the correct orientation on $K$ and $F_{K}^{\perp\kern-4.2pt\perp}$ is the closure of the other component of $\Sigma\setminus K$. Then $L=H_{1}(F_{K};\mathbb{Z})\subset H_{1}(\Sigma;\mathbb{Z})$ is a symplectic plane and $L^{\perp\kern-4.2pt\perp}=H_{1}(F_{K}^{\perp\kern-4.2pt\perp};\mathbb{Z})\subset H_{1}(\Sigma;\mathbb{Z})$.
Proof.
This is a statement about the intersection form on $H_{1}(\Sigma;\mathbb{Z})$. Up to applying an orientation preserving diffeomorphism, one may take $\Sigma$ to be one’s favorite genus $2$ surface, say $\Sigma_{\mathrm{std}}$ from Example 5.1, and one may take $F_{K}$ to be one’s favorite genus $1$ subsurface with one boundary component, say a closed small neighborhood of $\alpha_{1}\cup\beta_{1}$ from Example 5.1. We immediately have that $L$ is symplectic (e.g. by considering the basis $\alpha_{1}]$ and $\beta_{1}]$) and that $H_{1}(F_{K}^{\perp\kern-4.2pt\perp};\mathbb{Z})$ is its symplectic complement $L^{\perp\kern-4.2pt\perp}$ (e.g. by considering the basis $\alpha_{2}]$ and $\beta_{2}]$). ∎
Theorem 5.5.
Fix a closed oriented smooth genus $2$ surface $\Sigma\subseteq S^{3}$ and fix $D\equiv 1\pmod{4}$. Let $s_{1}$ and $s_{2}$ be in $\mathcal{Q}^{+}_{D}$. The following are equivalent:
  1. (a)
    There exists a non-contractible separating oriented simple closed curve $K\subseteq\Sigma$ such that the Seifert surfaces $F_{1}\coloneqq F_{K}$ and $F_{2}\coloneqq\overline{\Sigma\setminus(F_{K}\setminus K)}$ satisfy $s_{1}=s_{F_{1}}$ and $s_{2}=s_{F_{2}}$.
  2. (b)
    There exist $a,c\in\mathbb{Z}$ with 14ac=D1-4ac=D such that $ax^{2}+xy+cy^{2}]^{2}*s_{1}=s_{2}$.
Proof.
In this proof we identify $H_{1}(\Sigma;\mathbb{Z})$ with $\mathbb{Z}^{4}$ using the basis described in Lemma 5.2. In particular, via this identification $U_{\Sigma}$ is given by the matrix $U$ from (3.1).
Assume that a holds. Then $L=H_{1}(F_{1};\mathbb{Z})\subset H_{1}(\Sigma;\mathbb{Z})=\mathbb{Z}^{4}$ is a symplectic plane by Lemma 5.4. In particular, $s_{F_{1}}=[q_{L}]$. By Theorem 4.5, we have
[…]
Using Lemma 5.4 and Corollary 4.8, we find
[…]
Since $L$ is symplectic, $q_{a_{2}}$ is primitive by Lemma 3.5 and $\overline{[q_{a_{2}}]}=[q_{a_{2}}]^{-1}\in\mathcal{G}^{+}_{D}$. In fact, since $L$ is symplectic, $q_{a_{2}}=ax^{2}+xy+cy^{2}$ for some $a,c\in\mathbb{Z}$ with D=14acD=1-4ac. Hence, by acting with $q_{a_{2}}]^{-2}$ on $s_{1}$, we find
$ax^{2}+xy+cy^{2}]^{-2}*s_{1}=[q_{a_{2}}]^{-2}*\left(\overline{[q_{a_{1}}]}*[q_{a_{2}}]\right)=\overline{[q_{a_{1}}]}*\overline{[q_{a_{2}}]}=s_{2},$
so b follows from $ax^{2}+xy+cy^{2}]^{-1}=[cx^{2}+xy+ay^{2}]$ and relabeling.
Assume now that b holds. Set $q_{a_{2}}=ax^{2}-xy+cy^{2}$ and choose $q_{a_{1}}$ to be an IBQF in the class $q_{a_{2}}]^{-1}*\overline{s_{2}}=[q_{a_{2}}]*\overline{s_{1}}\in\mathcal{Q}^{+}_{D}$. The corresponding Klein vectors $a_{1}$ and $a_{2}$ define a symplectic plane $L$ of $\mathbb{Z}^{4}=H_{1}(\Sigma;\mathbb{Z})$ as in Proposition 3.2. Let $K\subset\Sigma$ be a non-contractible separating simple closed curve in $\Sigma$ such that $L=H_{1}(F_{K};\mathbb{Z})\subset H_{1}(\Sigma;\mathbb{Z})$, where $F_{K}$ is the closure of the component of $\Sigma\setminus K$ that induces the correct orientation on $K$; cf. by Lemma 5.3. Since $q_{L}]=s_{F_{K}}$ and, by Lemma 5.4, $s_{F_{2}}=s_{\overline{F_{K}^{\perp\kern-4.2pt\perp}}}=\overline{[q_{L^{\perp\kern-4.2pt\perp}}]}$, we have
[…]
and
[…]
as desired. ∎
Theorem 5.5 in particular establishes Theorem 1.1 from the introduction.
Proof of Theorem 1.1.
[…]
5.3. Obstruction to ambient isotopy in $\mathbf{B}^{4}$ and proof of Theorem 1.2
Let $F$ be a Seifert surface. We denote by $F^{\mathbf{B}^{4}}$ the neatly embedded compact surface in $\mathbf{B}^{4}$ obtained by pushing $F\setminus\partial F$ into the $4$-ball, i.e. we take $F^{\mathbf{B}^{4}}$ to be the time $1$ flow of a smooth vector field on $\mathbf{B}^{4}$ that is 0 on $\partial F$ and non-zero radially inward pointing on $S^{3}\setminus\partial F$. The surface $F^{\mathbf{B}^{4}}$ is well-defined up to smooth ambient isotopy relative to boundary.
Surfaces in $B^{4}$ arising as the result of pushing in a Seifert surface are in particular smooth submanifolds; however, the obstruction for differentiating two neatly embedded surfaces that we describe in what follows, can be formulated in the topological category. The later is worthwhile since we hope, in the future, to not only produce obstructions, but potentially use present ideas to provide classification results of pushed in Seifert surfaces up to topological ambient isotopy. However, in the topological category at many steps one ought to be careful to check things (e.g. well-definedness of double branched covers of ambiently isotopic surfaces), potentially invoking modern topological 4-manifold theory as e.g. outlined in [14]. We avoid considering these topological details, as in this article we only aim to differentiate pushed in Seifert surfaces, and only need the smooth setup.
For a neatly embedded compact surface $F$ in $\mathbf{B}^{4}$, we denote by $\Sigma(F)$ the oriented smooth (topological) $4$-manifold resulting from taking the double branched cover of $\mathbf{B}^{4}$ with branching set $F$. If two neatly embedded compact surfaces $F_{1}$ and $F_{2}$ in $\mathbf{B}^{4}$ can be mapped onto each other by an orientation preserving diffeomorphism (homeomorphism) of $\mathbf{B}^{4}$, then $\Sigma(F_{1})$ and $\Sigma(F_{2})$ are orientation preservingly diffeomorphic (homeomorphic). As, if two neatly embedded surfaces $F_{1}$ and $F_{2}$ in $\mathbf{B}^{4}$ are (topologically) ambiently isotopic, then, in particular, there exists an orientation preserving diffeomorphism (homeomorphism) $\phi\colon\mathbf{B}^{4}\to\mathbf{B}^{4}$ such that $\phi(F_{1})=F_{2}$, we find that the orientation preserving diffeomorphism (homeomorphism) type of the oriented smooth (topological) 4-manifold $\Sigma(F^{B^{4}})$ is well-defined.
For an oriented connected $4$-manifold $M$, we refer to the intersection form $H_{2}(M;\mathbb{Z})\times H_{2}(M;\mathbb{Z})\to\mathbb{Z}$ on the second integer homology group as the intersection form of $M$. (For completeness, recall that the intersection form is given by $H_{2}(M;\mathbb{Z})\times H_{2}(M;\mathbb{Z})\to\mathbb{Z};(A,B)\mapsto\left(\mathrm{PD}^{-1}(A)\cup\mathrm{PD}^{-1}(B)\right)([M])$, where $M]\in H_{4}(M;\mathbb{Z})$ denotes the fundamental class and $\mathrm{PD}\colon H^{2}(M;\mathbb{Z})\to H_{2}(M;\mathbb{Z})$ denotes Poincaré duality.) The intersection form of $\Sigma(F^{\mathbf{B}^{4}})$ is isomorphic to $Q_{F}$, meaning there exists an isometry (i.e. a group homomorphism that preserves the bilinear forms) between $H_{2}(\Sigma(F^{\mathbf{B}^{4}});\mathbb{Z})$ endowed with the intersection form and $H_{1}(F;\mathbb{Z})$ endowed with $Q_{F}$ [16, Theorem 3].
Lemma 5.6.
Let $F_{1}$ and $F_{2}$ be genus 1 Seifert surfaces for the same knot $K$. Then the following are equivalent:
  1. (a)
    $s_{F_{1}}$ is different from $s_{F_{2}}$ and $\overline{s_{F_{2}}}$, and
  2. (b)
    $\Sigma(F_{1}^{\mathbf{B}^{4}})$ and $\Sigma(F_{2}^{\mathbf{B}^{4}})$ have non-isometric intersection forms (and, in particular, are not orientation preservingly homeomorphic).
The key to Lemma 5.6 is the observation that, when $F$ is a genus $1$ Seifert surface of a knot $K$, the isomorphism type of $(H_{1}(F;\mathbb{Z}),Q_{F})$ is determined by the $\mathrm{GL}_{2}(\mathbb{Z})$-equivalence class of the IBQF $\frac{Q_{F}(a,a)}{2}x^{2}+Q_{F}(a,b)xy+\frac{Q_{F}(b,b)}{2}y^{2}$, where $(a,b)$ is any basis for $H_{1}(F;\mathbb{Z})$.
Proof of Lemma 5.6.
Using that the intersection form on $H_{2}(\Sigma(F_{i}^{B^{4}});\mathbb{Z})$ is isomorphic to $Q_{F_{i}}$ on $H_{1}(F_{i};\mathbb{Z})$, we have that b holds if and only if there does not exist $M\in\mathrm{GL}_{2}(\mathbb{Z})$ such that $q_{F_{1}}((x,y)M)=q_{F_{2}}$. Here, we have
$q_{F_{i}}$ $\coloneqq U_{F_{i}}(a_{i},a_{i})x^{2}+(U_{F_{i}}(a_{i},b_{i})+U_{F_{i}}(b_{i},a_{i}))xy+U_{F_{i}}(b_{i},b_{i})y^{2}$
$=\frac{1}{2}\left(Q_{F}(a_{i},a_{i})x^{2}+2Q_{F_{i}}(a_{i},b_{i})xy+Q_{F}(b_{i},b_{i})y^{2}\right)$
for $(a_{i},b_{i})$ some symplectic basis of $H_{1}(F_{i};\mathbb{Z})$. The above is equivalent to $s_{F_{1}}\notin\{s_{F_{2}},\overline{s_{F_{2}}}\}$, that is to a, as desired. ∎
Note that in the following theorem, which is the more precise version of Theorem 1.2, the case of primitive $s$ corresponds to $s$ with content $1$.
Theorem 5.7.
Fix an integer $D\equiv 1\pmod{4}$. The following are equivalent.
  1. (a)
    The subgroup $\mathcal{S}^{+}_{D}$ of $\mathcal{G}^{+}_{D}$ is non-trivial, i.e. there exist $a,c\in\mathbb{Z}$ with 14ac=D1-4ac=D such that $ax^{2}+xy+cy^{2}]^{2}\neq 1$.
  2. (b)
    There exist a knot with determinant $D$ and Seifert surfaces $F_{1}$ and $F_{2}$ such that $\Sigma(F_{1}^{\mathbf{B}^{4}})$ and $\Sigma(F_{2}^{\mathbf{B}^{4}})$ have non-isometric intersection forms (and, in particular, are not orientation preservingly homeomorphic).
Furthermore, for every $m\geq 1$ with $m^{2}\mid D$, the following are equivalent.
  1. (1)
    The image of the subgroup $\mathcal{S}_{D}^{+}$ under the map $\mathcal{G}_{D}^{+}\to\mathcal{G}_{D/m^{2}}^{+}$ (cf. Corollary 4.4) has order at least $3$.
  2. (2)
    For all $s\in\mathcal{Q}_{D}^{+}$ with content $m$, there exist Seifert surfaces $F_{1}$ and $F_{2}$ such that $s_{F_{1}}=s$ and $\Sigma(F_{1}^{\mathbf{B}^{4}})$ and $\Sigma(F_{2}^{\mathbf{B}^{4}})$ have non-isometric intersection forms.
In fact, as will be clear from the proof, the Seifert surfaces in b can be chosen to be disjoint.
Proof.
To see a implies b, assume a holds and let $t=[ax^{2}+xy+cy^{2}]\in\mathcal{G}^{+}_{D}$ with 14ac=D1-4ac=D such that $t^{2}\neq 1$. We set $s_{1}=t^{2}$ and $s_{2}=1$. By Theorem 5.5, there exists a knot $K$ with disjoint Seifert surfaces $F_{1}$ and $F_{2}$ such that $s_{F_{1}}=s_{1}$ and $s_{F_{2}}=s_{2}$; in particular, $K$ has determinant $D$. By Lemma 5.6, $\Sigma(F_{1}^{\mathbf{B}^{4}})$ and $\Sigma(F_{2}^{\mathbf{B}^{4}})$ have non-isometric intersection forms.
We give two proofs to see b implies a. Let $F_{1}$ and $F_{2}$ be genus 1 Seifert surfaces for the same knot $K$. By [26], there exists a sequence of genus 1 Seifert surfaces $G_{1},G_{2},\cdots,G_{n}$ such that $G_{i}\cap G_{i+1}=K$ for all $1\leq i<n$ and $G_{1}=F_{1}$ and $G_{n}=F_{2}$. Hence, by applying Theorem 5.5 to $\Sigma\coloneqq G_{i}\cup\overline{G_{i+1}}$, we find $t_{i}=[a_{i}x^{2}+xy+c_{i}y^{2}]$ such that $t_{i}^{2}*s_{G_{i}}=s_{G_{i+1}}$. Therefore, $(\prod_{i=1}^{n-1}t_{i}^{2})*s_{F_{1}}=s_{F_{2}}$. Now assume $F_{1}$ and $F_{2}$ are as described in b. In particular, $s_{F_{1}}\neq s_{F_{2}}$, hence $t_{i}^{2}\neq 1$ for at least one $i$, and by setting $a=a_{i}$ and $c=c_{i}$ for this $i$, we have proven a. Alternatively, this can be proved via results of the third author from [23, §2] on S-equivalence of Seifert pairings of genus 1 knots; see Remark 5.11 in Section 5.4, where S-equivalence and the aforementioned results by the third author are discussed.
To see the last part of the theorem, we note that (1) is equivalent to the following. There exists $t_{1},t_{2}\in\mathcal{G}_{D}^{+}$ of the form $t_{i}=[a_{i}x^{2}+xy+c_{i}y^{2}]$ where $D=1-4a_{i}c_{i}$ such that the images of $t_{1}^{2}$ and $t_{2}^{2}$ are distinct and non-trivial under the map $\mathcal{G}_{D}^{+}\to\mathcal{G}_{D/m^{2}}^{+}$ (cf. Corollary 4.4).
Assume first $t_{1},t_{2}\in\mathcal{G}_{D}^{+}$ are as in the equivalent formulation of (1) above. Set $s_{1}=s$ and $s_{2}=t_{1}^{2}*s$ in a first attempt. Then $s_{1}\neq s_{2}$ (because $\tilde{t}_{1}$ is not $2$-torsion) and $s_{1}=\overline{s_{2}}$ if and only if $(\frac{1}{m}s)^{2}=\tilde{t}_{1}^{-2}$. If the latter occurs, we replace $t_{1}$ with $t_{2}$ and note that $(\frac{1}{m}s)^{2}=\tilde{t}_{1}^{2}\neq\tilde{t}_{2}^{2}$ to ensure $s_{1}\neq\overline{s_{2}}$. By construction, $s_{1}\notin\{s_{2},\overline{s_{2}}\}$ and so $\Sigma(F_{1}^{\mathbf{B}^{4}})$ and $\Sigma(F_{2}^{\mathbf{B}^{4}})$ have non-isometric intersection forms as desired.
Assume lastly that Item (2) holds, and apply it for $s_{F_{1}}=m\cdot 1_{D/m^{2}}$. By Theorem 5.5 we find $t_{1}$ with $t_{1}^{2}\in\mathcal{S}_{D}^{+}$ such that
$t_{1}^{2}*(m\cdot 1_{D/m^{2}})=m(\tilde{t}_{1}^{2}*1_{D/m^{2}})=s_{F_{2}}\neq s_{F_{1}}=m\cdot 1_{D/m^{2}},$
where we used (4.3) in the last equality. It follows that $\tilde{t}_{1}^{2}\neq 1_{D/m^{2}}$. We now apply Item (2) in the same way, but with $s_{1}:=s_{F_{1}}=m\tilde{t}_{1}$, which has content $m$. By Theorem 5.5 we find $t_{2}$ with $t_{2}^{2}\in\mathcal{S}_{D}^{+}$ such that
$s_{2}:=s_{F_{2}}=t_{2}^{2}*s_{1}=t_{2}^{2}*(m\tilde{t}_{1})=m(\tilde{t}_{2}^{2}*\tilde{t}_{1})$
and with $s_{1}\notin\left\{s_{2},\overline{s_{2}}\right\}$. Equivalently,
$m\tilde{t}_{1}\neq m(\tilde{t}_{2}^{2}*\tilde{t}_{1}),\quad m\tilde{t}_{1}\neq\overline{s_{2}}=m(\tilde{t}_{2}^{-2}*\tilde{t}_{1}^{-1}).$
This gives $t_{1},t_{2}\in S_{D}^{+}$ with $\tilde{t_{2}}^{2}\neq 1_{D/m^{2}}$ and with $\tilde{t_{1}}^{2}\neq\tilde{t_{2}}^{-2}$. Together with $\tilde{t_{1}}^{2}\neq 1_{D/m^{2}}$ that was established above, Item (1) follows. ∎
This result in particular gives the less technical Theorem 1.2 from the introduction, using that $s\in\mathcal{Q}^{+}_{D}$ is primitive exactly means that $s$ has content $1$.
Proof of Theorem 1.2.
We note that by Theorem 5.7, there exists a knot with determinant $D$ and disjoint Seifert surfaces $F_{1}$ and $F_{2}$ such that $\Sigma(F_{1}^{\mathbf{B}^{4}})$ and $\Sigma(F_{2}^{\mathbf{B}^{4}})$ are not orientation preservingly homeomorphic. Hence, there cannot exist an orientation preserving self-homeomorphism $\phi\colon B^{4}\to B^{4}$ with $\phi(F_{1})=F_{2}$; i.e. $F_{1}^{\mathbf{B}^{4}}$ and $F_{2}^{\mathbf{B}^{4}}$ are not ambiently isotopic.
For the second part of the theorem, assume there exists $t=[ax^{2}+xy+cy^{2}]\in\mathcal{G}^{+}_{D}$ such that $t^{4}\neq 1$, then in particular $t^{2}$ is an element of order at least 3 and the statement follows from the furthermore of Theorem 5.7: we find a knot with surfaces $F_{1}$ and $F_{2}$ such that $s_{F_{1}}$ is any fixed $s\in\mathcal{G}^{+}_{D}$ (i.e. $s\in\mathcal{Q}^{+}_{D}$ with content $1$). ∎
5.4. S-equivalence
We comment further on the condition in Theorem 1.1 respectively Theorem 5.5, which states that IBQFs $s_{1}$ and $s_{2}$ can be realized as the Seifert forms of a disjoint genus $1$ pair if and only if there exists a class $t=[ax^{2}+xy+cy^{2}]$ such that $t^{2}*s_{1}=s_{2}$.
After discovering Theorem 1.1 through an algebraic calculation, the authors realize that the necessary and sufficient condition fits with the following prior result by the third author. For a fixed discriminant $D\equiv 1\pmod{4}$, recall that we denote by $\mathcal{S}^{+}_{D}$ the subgroup of $\mathcal{G}^{+}_{D}$ generated by elements of the form $ax^{2}+xy+cy^{2}]^{2}$. The third author has described and investigated the notion of S-equivalence for Seifert forms of Seifert surfaces of genus 1 for knots using this subgroup [23, §2]. In doing so, she recovered some results that were found much earlier by Dean Bandes in his Ph.D. thesis [2]. In the notation of the present paper, the relevant result is the following.
Theorem 5.8 ([2, Thm. 14.1], [23, Prop. 2.9]).
Let $s_{1}$ and $s_{2}$ be in $\mathcal{Q}^{+}_{D}$. Then $s_{1}$ and $s_{2}$ are S-equivalent if and only if there exists an element $s\in\mathcal{S}^{+}_{D}$ such that $s*s_{1}=s_{2}$. In other words, $s_{1}$ and $s_{2}$ are S-equivalent if and only if there exists a finite number of $t_{1},\dots,t_{n}$ of the form $t_{i}=[a_{i}x^{2}+xy+c_{i}y^{2}]$ such that $(\prod_{i=1}^{n}t_{i}^{2})*s_{1}=s_{2}$.
Theorem 5.8 is a translation of [23, Prop. 2.9] into the language of binary quadratic forms via the correspondence described in [23, §2]. Alternatively, it is a translation of [2, Thm. 14.1] into the language of binary quadratic forms via the correspondence described in [2, Thm. 9.1]. (The strict similarity classes of modules defined in [2] are essentially narrow ideal classes.)
While we abstain from defining the notion of S-equivalence beyond giving the above equivalent formulation, we note that the notion is of interest and arises naturally in knot theory due to the following.
Lemma 5.9 ([20, Theorem 8.4]).
All Seifert forms that arise from Seifert surfaces with boundary a fixed knot are S-equivalent.
To clarify the connection between our result and the earlier work of the third author, we describe how to use Theorem 1.1 respectively Theorem 5.5 to recover the following statement, which is the easier half of Theorem 5.8.
Corollary 5.10.
Fix $D\equiv 1\pmod{4}$. Let $s_{1}$ and $s_{2}$ be in $\mathcal{Q}^{+}_{D}$. If there exists an $s\in\mathcal{S}^{+}_{D}$ such that $s*s_{1}=s_{2}$, then $s_{1}$ and $s_{2}$ are S-equivalent.
Proof.
By the definition of $\mathcal{S}^{+}_{D}$, there exist $t_{1},\dots t_{n}$ of the form $t_{i}=[a_{i}x^{2}+xy+c_{i}y^{2}]$ such that $s=\prod_{i=1}^{n}t_{i}^{2}$. By Theorem 5.5, for all $1\leq k\leq n$, there exists a knot $K_{k}$ with a pair of genus $1$ Seifert surfaces $F_{k}$ and $F_{k}^{\prime}$ (that are disjoint in their interior) such that $\prod_{i=1}^{k-1}t_{i}^{2}*s_{1}=s_{F_{k}}$ and $\prod_{i=1}^{k}t_{i}^{2}*s_{1}=s_{F_{k}^{\prime}}$. In particular, for all $1\leq k\leq n-1$, $\prod_{i=1}^{k-1}t_{i}^{2}*s_{1}=s_{F_{k}}$ and $\prod_{i=1}^{k}t_{i}^{2}*s_{1}=s_{F_{k}^{\prime}}$ are S-equivalent by Lemma 5.9. Since S-equivalence is an equivalence relation and as such is transitive, we find $s_{1}$ and $s_{2}=\prod_{i=1}^{n}t_{i}^{2}*s_{1}$ are S-equivalent. ∎
Note that in the above proof, we did not establish that $s_{1}$ and $s_{2}$ can be realized as the Seifert forms of Seifert surfaces for the same knot. However, given Theorem 5.5, which for the special case of $s*s_{1}=s_{2}$ with $s=[ax^{2}+xy+cy^{2}]^{2}$ not only establishes S-equivalence but in fact realization as Seifert forms for Seifert surfaces for the same knot, we conjecture the following. For all $D\equiv 1\pmod{4}$, if $s\in\mathcal{S}^{+}_{D}$ and $s_{1},s_{2}\in\mathcal{Q}^{+}_{D}$ satisfy $s*s_{1}=s_{2}$, then there exist genus $1$ Seifert surfaces $F_{1}$ and $F_{2}$ with boundary the same knot such that $s_{1}=s_{F_{1}}$ and $s_{2}=s_{F_{2}}$; see Problem 7.7.
Remark 5.11.
We provide a proof of b implies a of Theorem 5.7 using Theorem 5.8 in place of making use of [26]. Assume $F_{1}$ and $F_{2}$ are Seifert surfaces as described in b. By Lemma 5.9, $s_{F_{1}}$ and $s_{F_{2}}$ are S-equivalent; hence, by Theorem 5.8, there exist $t_{1},\dots,t_{n}$ of the form $t_{i}=[a_{i}x^{2}+xy+c_{i}y^{2}]$ such that $(\prod_{i=1}^{n}t_{i}^{2})*s_{F_{1}}=s_{F_{2}}$. Now, as in the proof of Theorem 5.7, since $s_{F_{1}}\neq s_{F_{2}}$, $t_{i}^{2}\neq 1$ for at least one $i$.
We end this paragraph with more philosophical comments. Firstly, in hindsight, in light of the third author’s earlier results, the condition we found in Theorem 1.1 respectively Theorem 5.5 make sense, as they are a refinement of what needs to hold for S-equivalent $s_{1}$ and $s_{2}$. However, this is not how we discovered them. They arose by direct calculation using Theorem 4.5, we only later noticed the connection. Secondly, the strength of Theorem 5.5 (certainly with view of the application to pushed in Seifert surfaces) lies in the fact the IBQFs are realized by Seifert surfaces with the same knot as its boundary. Instead, the third author’s prior result is an algebraic statement and does not give realization for knots (however it does imply a realization result in the analogous high-dimensional setting of simple $(4a+1)$-knots, where two knots with S-equivalent Seifert pairings necessarily bound isotopic knots). The natural analog of the third author’s result (Theorem 5.8) would be the positive resolution of the conjecture made in the above paragraph. As an immediate application, one could find many further examples of forms that are realized by Seifert surfaces with the same boundary and hence could use that to provide many more examples of Seifert surfaces that are non-isotopic even when pushed into the $4$-ball.
5.5. Realization for fixed knots
The reader may notice that in our main 3-dimensional results (i.e. Theorem 1.1 and the more detailed Theorem 5.5), we make no claim about the knot that arises. A stronger claim would be the following. Given a genus 1 Seifert surface $F$ for a knot $K$ such that $s_{F}=s_{1}$ and $s_{2}$ with $ax^{2}+xy+cy^{2}]^{2}*s_{1}=s_{2}$ for some $a,c\in\mathbb{Z}$, there exist Seifert surfaces $F_{1}$ and $F_{2}$ for the knot $K$ such that $s_{1}=s_{F_{1}}$ and $s_{2}=s_{F_{2}}$. Or even stronger, one might hope to find those $F_{1}$ and $F_{2}$ with the additional restriction that $F_{1}$ and $F_{2}$ are disjoint away from the boundary and that $F_{1}=F$. In fact, if either of the above were true, in the proof of Corollary 5.10, we could choose all $K_{i}$ to be the same knot and would obtain a much stronger result, namely that, if $s*s_{1}=s_{2}$ for some $s\in\mathcal{S}^{+}_{D}$, then $s_{1}$ and $s_{2}$ are realized as the Seifert forms of genus 1 Seifert surfaces with boundary the same knot; compare with the conjecture in the above subsection and with Problem 7.7.
However, such stronger claims cannot hold in general. In fact, even for the case of D=23D=-23, there is an example of a knot for which such a result cannot be obtained.
Example 5.12.
The knot $K=9_{5}$ has exactly two non-isotopic Seifert surfaces $F$ and $F^{\prime}$. These surfaces have associated Seifert forms $3x^{2}+xy+2y^{2}]$ and $3x^{2}-xy+2y^{2}]$ respectively. (In fact, $F$ and $F^{\prime}$ become smoothly ambiently isotopic relative boundary when pushed into the $4$-ball.) We elaborate. To see two distinct surfaces, we take $F$ and $F^{\prime}$ as the result obtained by plumbing together a $3$-times twisted unknotted band to a $2$-times twisted unknotted band as indicated in Figure 2.
$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\beta$}
$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\beta$}
$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\alpha$}
$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\alpha$}
Figure 2. The two Seifert surfaces (gray) for $K$. The simple closed curves $\alpha$ and $\beta$ (red) are such that $([\alpha],[\beta])$ forms a symplectic basis of the first homology of the surfaces.
They are readily seen to have Seifert forms given by $\big(\begin{smallmatrix}2&1\\ 0&3\end{smallmatrix}\big)$ and $\big(\begin{smallmatrix}2&0\\ -1&3\end{smallmatrix}\big)$ with respect the basis $\alpha]$ and $\beta]$ of first homology indicated in Figure 2. The Seifert surfaces are non-isotopic in $S^{3}$ since $2x^{2}+xy+3y^{2}]\neq[2x^{2}-xy+3y^{2}]\in\mathcal{G}_{-23}^{+}$. In fact, the surfaces become isotopic if the orientation on one of them is reversed, which is reflected in the fact that $2x^{2}+xy+3y^{2}]=\overline{[2x^{2}-xy+3y^{2}]}$. There are no other genus 1 Seifert surfaces up to isotopy by [18, Theorem A].
As a result, we have a knot $K$ with a genus 1 Seifert surface $F$, such that there does not exist a Seifert surface $F_{2}$ of genus 1 with $s_{F_{2}}=1\in\mathcal{G}_{-23}^{+}$, even though there are $a$ and $c$ such that $ax^{2}+xy+cy^{2}]^{2}*s_{1}=s_{2}$, where $s_{1}=s_{F}=[3x^{2}+xy+2y^{2}]$ and $s_{2}=1$.
6. Examples
In this section we prove Corollary 1.3 and give a variety of examples to illustrate the main results. The explicit calculations in the examples below can e.g. be done using standard descriptions of Gauss composition [15], [7, Chapt. 14]. We implemented and verified most of them using SageMath.
6.1. Negative discriminants
We begin the discussion with proving Corollary 1.3. The following is the main observation.
Lemma 6.1.
Suppose that $D$ is a negative discriminant with $D\equiv 1\,\mathrm{mod}\,4$. Then $ax^{2}+xy+cy^{2}]^{2}=\mathrm{id}$ for all $a,c\in\mathbb{Z}$ with D=14acD=1-4ac if and only if $D$ is of the form $D=1-4p^{k}$ for $k\in\{0,1,2\}$ and $p$ a prime.
Proof.
We first claim that the square of any class $ax^{2}+xy+cy^{2}]$ is given by
(6.1) $[ax^{2}+xy+cy^{2}]^{2}=[a^{2}x^{2}+(1-2ac)xy+c^{2}y^{2}].$
This can be checked directly by Gauss composition. Alternatively, and in the spirit of the current article, let us recall that the square class $ax^{2}+xy+cy^{2}]^{2}$ is exactly minus the class of the quadratic form on the oriented plane $\langle 1,A\rangle^{\perp}$ where $A=\big(\begin{smallmatrix}1&-2a\\ 2c&-1\end{smallmatrix}\big)$ (cf. Corollary 4.8). An oriented (integer) basis of $\langle 1,A\rangle^{\perp}=L_{A,-A}$ is easily computed to be $\big(\begin{smallmatrix}a&0\\ 1&-a\end{smallmatrix}\big)$ and $\big(\begin{smallmatrix}-c&1\\ 0&c\end{smallmatrix}\big)$ which implies (6.1).
We may assume $a,c>0$. As the trivial class $x^{2}+xy+acy^{2}$ primitively represents no integers less than $ac$ except for $1$, the square class (6.1) is non-trivial unless $a=1$, $c=1$, or $a=c$ (in which case it is trivial). Turning to the statement of Lemma 6.1, $D$ may be written as D=14acD=1-4ac for $a,c>1$ with $a\neq c$ if and only if it is not of the form $D=1-4p^{k}$ for $k\in\{0,1,2\}$ and $p$ a prime. In this case exactly, we can find a non-trivial class as in (6.1) and hence the proposition follows. ∎
Proof of Corollary 1.3.
This follows directly from Lemma 6.1 and Theorem 5.7. ∎
In the following example we discuss the discriminant D=23D=-23 in full detail. In particular, this connects to [17] and extends upon Example 1.4.
Example 6.2.
The oriented class group for discriminant 23-23 has representatives
$\mathcal{G}_{D}^{+}=\{[x^{2}+xy+6y^{2}],[-x^{2}-xy-6y^{2}],[2x^{2}\pm xy+3y^{2}],[-2x^{2}\pm xy-3y^{2}]\}.$
As a group of order $6$, it is isomorphic to $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/3\mathbb{Z}$. Recall from Example 1.4 that by Theorem 1.1, two Seifert forms $s_{1},s_{2}\in\mathcal{G}^{+}_{-23}$ arise from a disjoint genus 1 pair of Seifert surfaces if and only if $2x^{2}\pm xy+3y^{2}]*s_{1}=s_{2}$ or $s_{1}=s_{2}$. The unordered pairs of such $s_{1}$ and $s_{2}$ are, in addition to the six with $s_{1}=s_{2}$, the following:
$\{1,[2x^{2}+xy+3y^{2}]\},\{1,[2x^{2}-xy+3y^{2}]\},\{[-x^{2}-xy-6y^{2}],[-2x^{2}-xy-3y^{2}]\},$
$\{[-x^{2}-xy-6y^{2}],[-2x^{2}+xy-3y^{2}]\},\{[2x^{2}+xy+3y^{2}],[2x^{2}-xy+3y^{2}]\},$
$\text{ and }\quad\{[-2x^{2}-xy-3y^{2}],[-2x^{2}+xy-3y^{2}]\}.$
For a pair of Seifert surfaces of the same knot that realize $s_{1}$ and $s_{2}$, we have by Lemma 5.6 that the surfaces are not ambient isotopic in $B^{4}$ whenever $s_{1}^{-1}=\overline{s_{1}}\neq s_{2}\neq s_{1}$. Of the above pairs with $s_{1}\neq s_{2}$, only the last two satisfy $s_{1}^{-1}=s_{2}$. Hence, we have that the other $4$ pairs arise from Seifert surfaces as asked for by Livingston. The example chosen in [17, Proof of Theorem 4.1] is the fourth of the above pairs: $x^{2}-xy-6y^{2}]$ and $2x^{2}+xy-3y^{2}]=[-6x^{2}-5xy-2y^{2}]$. The other three pairs are obtained from this one by taking mirror image (switches from negative definite to positive definite) and reversing orientation of the Seifert surfaces (corresponds to taking inverses).
Observe that in the above example a pair $s_{1},s_{2}$ arises from a disjoint genus one pair of Seifert surfaces if and only if $s_{1},s_{2}$ are S-equivalent i.e. there exists $s\in\mathcal{S}_{D}^{+}$ with $s*s_{1}=s_{2}$. This is not generally true, as the example of D=71D=-71 shows.
Example 6.3.
Consider D=71D=-71 and note that $71=1-4\cdot 2\cdot 9$. The non-trivial classes of discriminant 71-71 of the form $ax^{2}+xy+cy^{2}]^{2}$, $ac=18$, are $3x^{2}\pm xy+6y^{2}]^{2}=[2x^{2}\mp xy+9y^{2}]$ and $2x^{2}\pm xy+9y^{2}]^{2}=[4x^{2}\mp 3xy+5y^{2}]$. The oriented class group $\mathcal{G}_{-71}^{+}$ is a group of order $14$ where the (non-oriented) class group $\mathcal{G}_{D}\cong\mathbb{Z}/7\mathbb{Z}$ is
$\mathcal{G}_{D}=\{[x^{2}+xy+18y^{2}],[2x^{2}\pm xy+9y^{2}],[3x^{2}\pm xy+6y^{2}],[4x^{2}\pm 3xy+5y^{2}]\}.$
The cyclic group $\mathbb{Z}/7\mathbb{Z}$ is generated by any non-trivial element and hence $\mathcal{S}_{-71}^{+}=(\mathcal{G}_{-71}^{+})^{2}$. On the other hand, it is clear that (for example) the classes $s_{1}=[x^{2}+xy+18y^{2}],s_{2}=[3x^{2}+xy+6y^{2}]$ are not related by multiplication with a class of the form $ax^{2}+xy+cy^{2}]^{2}$. In particular, our results do not answer whether or not there exist a knot $K$ and two genus one Seifert surfaces $F_{1},F_{2}$ with boundary $K$ such that $s_{F_{1}}=s_{1}$ and $s_{F_{2}}=s_{2}$. However, Theorem 1.1 shows that such $F_{1},F_{2}$, if they exist, cannot be disjoint in their interior. We also refer to Problem 7.7 for a general formulation.
Example 6.4.
Fehér [13] constructs an infinite set of examples of genus one disjoint pairs of Seifert surfaces and shows that they are not ambiently isotopic in the $4$-ball. In the following, we discuss their work in the context of the current article. Let $p,q>1$ be coprime integers, $n\geq 1$ (which will later be chosen large), and $k\in\mathbb{Z}$. Choose $r,s\in\mathbb{Z}$ with psqr=1ps-qr=1. Then there exists a symplectic plane $L$ such that
$[q_{L}]$ $=[pqx^{2}+(1-2kp)xy+ny^{2}],$
$[q_{L^{\perp\kern-4.2pt\perp}}]$ $=[pqx^{2}+(2kp-1-2qr)xy+(rs-2kr+n)y^{2}].$
Fehér establishes this by drawing disjoint genus one pairs of Seifert surfaces realizing these forms. In the context of this article, one may verify that the plane $L=L_{a_{1},a_{2}}$ for
$a_{1}=\begin{pmatrix}-1+2kp&2q\\ -2np&1-2kp\end{pmatrix},\ a_{2}=\begin{pmatrix}1&2p\\ 2(k^{2}p-k-nq)&-1\end{pmatrix}$.
gives rise to the desired forms.
Fehér [13] verifies by direct elementary means that $q_{L}]\neq[q_{L^{\perp\kern-4.2pt\perp}}],[\overline{q_{L^{\perp\kern-4.2pt\perp}}}]$ whenever $2kp\not\equiv 1\,\mathrm{mod}\,q$ and
$n\geq\frac{k(kp-1)}{q}+\frac{pq}{12}-\frac{1}{6}-\frac{1}{2pq}.$
In particular, this shows that the corresponding disjoint genus one pairs of Seifert surfaces remain non-isotopic when pushed into the $4$-ball. Corollary 4.9 implies that $q_{L}]\neq[q_{L^{\perp\kern-4.2pt\perp}}],[\overline{q_{L^{\perp\kern-4.2pt\perp}}}]$ if and only if
(6.2) $[q_{a_{1}}]\neq[\overline{q_{a_{1}}}],\quad[q_{a_{2}}]^{2}\not\in\mathrm{Stab}([q_{a_{1}}]).$
As in Lemma 6.1 one may verify that this is, for instance, the case if $2kp\not\equiv 1\,\mathrm{mod}\,q$, $\gcd(1-2kp,q,n)=1$, and
$n>\frac{k(kp-1)+p}{q}+\frac{1}{4pq}.$
Moreover, for specific values of $p,q,k,n$ the conditions in (6.2) can be quickly verified using reduction theory.
6.2. Positive discriminants
Generally, determining class groups for positive discriminants is (much) harder than for negative discriminants (see e.g. Gauss’s class number problem).
Example 6.5.
The discriminant $D=145$ is the smallest positive squarefree discriminant for which there exists a non-trivial special class $ax^{2}\pm xy+cy^{2}]^{2}$ with ac=(D1)/4=36ac=-(D-1)/4=-36. Such a non-trivial class is given by $2x^{2}+xy-18y^{2}]^{2}=[6x^{2}+5xy-5y^{2}]$. We note that the oriented class group in this case is isomorphic to $\mathbb{Z}/4\mathbb{Z}$ with generator $3x^{2}+7xy-8y^{2}]$.
Example 6.6.
The discriminant $D=905$ is a positive and squarefree discriminant for which there exist a class $s\in(\mathcal{G}_{D}^{+})^{2}=(\mathcal{G}_{D})^{2}$ which is not of the form $ax^{2}\pm xy+cy^{2}]^{2}$ for any $a,c$ with 14ac=D1-4ac=D. In fact, a calculation verifies that $\mathcal{G}_{D}^{+}\simeq\mathbb{Z}/8\mathbb{Z}$ so that $(\mathcal{G}_{D}^{+})^{2}\simeq\mathbb{Z}/4\mathbb{Z}$. As $(905-1)/4=226=2\cdot 113$, there exist at most $3$ classes of the form $ax^{2}\pm xy+cy^{2}]^{2}$ which verifies the above claim. This situation is analogous to Example 6.3.
6.3. Square discriminants
6.3.1. Reduction theory and Gauss composition for square discriminants
Many treatments of the theory of binary quadratic forms exclude the case where the determinant $D$ is a square number. Although this case is in some regards arithmetically less interesting than the others, as observed in [3], Gauss composition still makes sense in this context. In fact, as we will see below, the group $\mathcal{G}^{+}_{D}$ admits an explicit description when $D=N^{2}$: it is isomorphic to the multiplicative group modulo $N$.
From a topological viewpoint, the case $D=N^{2}$ is exactly the case when the Alexander polynomial factors over the integers as
(6.3) $mt+(1-2m)+mt^{-1}=(kt-(k+1))(kt^{-1}-(k+1)).$
where D=14mD=1-4m and $N=2k+1>0$. These polynomials have a topological meaning – they are exactly those which appear as Alexander polynomials of genus 1 slice knots. A genus 1 knot with Alexander polynomial of the form (6.3) does not have to be slice; however, it is algebraically slice, that is, there is a $1$-dimensional isotropic subspace for the Seifert form of any genus $1$ Seifert surface of $K$. The latter in particular follows from Proposition 6.7.
For lack of (explicit) reference, we now write down the details of reduction theory and Gauss composition for the case of odd square discriminants $D=N^{2}$, $N=2k+1>0$.
Proposition 6.7 (Reduction Theory for Square Discriminant).
Any quadratic form of discriminant $D=N^{2}$ is $\operatorname{SL}_{2}(\mathbb{Z})$-equivalent to a unique one of the forms
$Q_{N,a}=ax^{2}+Nxy+0y^{2}.$
as $a$ runs through a complete set of representatives of the residue classes modulo $N$.
Proof.
Since any quadratic form is a positive scalar multiple of a primitive form, it suffices to show this for primitive quadratic forms.
Let $Q$ be a primitive quadratic form of discriminant $D=N^{2}$. Since the discriminant is square, $Q$ factors over the rationals, and so also over the integers by Gauss’s lemma. Write $Q(x,y)=(\alpha x+\gamma y)(\beta x+\delta y)$ for integers $\alpha,\beta,\gamma,\delta$. One verifies that for $g\in\operatorname{SL}_{2}(\mathbb{Z})$ we have
(6.4) $g.Q(x,y)=(\alpha^{\prime}x+\gamma^{\prime}y)(\beta^{\prime}x+\delta^{\prime}y)$
where $\left(\begin{smallmatrix}\alpha^{\prime}&\beta^{\prime}\\ \gamma^{\prime}&\delta^{\prime}\end{smallmatrix}\right)=g\cdot\left(\begin{smallmatrix}\alpha&\beta\\ \gamma&\delta\end{smallmatrix}\right)$. One can compute $N^{2}=\operatorname{disc}Q=\det\left(\begin{smallmatrix}\alpha&\beta\\ \gamma&\delta\end{smallmatrix}\right)^{2}$. Swapping the columns of the matrix flips the sign of the determinant, so we can choose our labeling of the factors so that $\det\left(\begin{smallmatrix}\alpha&\beta\\ \gamma&\delta\end{smallmatrix}\right)=N$.
By Gauss’s lemma, primitivity of $Q$ is equivalent to primitivity of the vectors $(\alpha,\gamma)$ and $(\beta,\delta)$. Thus, we may choose a matrix $g\in\operatorname{SL}_{2}(\mathbb{Z})$ such that $g\left(\begin{smallmatrix}\alpha\\ \gamma\end{smallmatrix}\right)=\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)$. By the above
(6.5) $g\begin{pmatrix}\alpha&\beta\\ \gamma&\delta\end{pmatrix}=\begin{pmatrix}1&a\\ 0&N\end{pmatrix}$
It follows from (6.4) that
(6.6) $g.Q(x,y)=x(ax+Ny)=ax^{2}+Nxy+0y^{2}.$
Note however that $g$ is only well-defined up to left multiplication by the stabilizer $\left(\begin{smallmatrix}1&k\\ 0&1\end{smallmatrix}\right)$ of the vector $\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)$. This action sends $ax^{2}+Nxy+0y^{2}$ to $(a+Nk)x^{2}+Nxy+0y^{2}$, so $a$ is only well-defined modulo $N$. We can choose $g$ such that $a$ lands in a given set of representatives of the primitive residue classes modulo $N$.
To obtain uniqueness, let $0\leq a,a^{\prime}<N$ be representatives of primitive residue classes modulo $N$ and suppose that $g\in\operatorname{SL}_{2}(\mathbb{Z})$ is such that $g.Q_{N,a}=Q_{N,a^{\prime}}$. By (6.4) we have either
[…]
In the former case, this implies $a=a^{\prime}$. In the latter case, $g=\big(\begin{smallmatrix}a^{\prime}&(-aa^{\prime}+1)/N\\ N&-a\end{smallmatrix}\big)$ which has determinant 1-1; hence this case does not occur. This proves the lemma.∎
Proposition 6.8 (Gauss Composition for Square Discriminant).
When $D=N^{2}$ is a square discriminant, the group of $\operatorname{SL}_{2}(\mathbb{Z})$-equivalence classes of primitive binary quadratic forms under Gauss composition is canonically isomorphic to the multiplicative group modulo $N$ via the following map
(6.7) $\phi_{N}:(\mathbb{Z}/N\mathbb{Z})^{\times}\overset{\sim}{\longrightarrow}\mathcal{G}^{+}_{N^{2}},\quad a\bmod N\mapsto[Q_{N,a}]=[ax^{2}+Nxy+0y^{2}].$
Proof.
The map $\phi_{N}$ is well-defined because $ax^{2}+Nxy]$ is $\operatorname{SL}_{2}(\mathbb{Z})$-equivalent to $(a+kN)x^{2}+Nxy]$ for all integers $k$. Proposition 6.7 then says precisely that $\phi_{N}$ is a bijection. Applying the Dirichlet composition law applied to the two concordant forms $a_{1}x^{2}+Nxy+0y^{2}$ and $a_{2}x^{2}+Nxy+0y^{2}$ (cf. [7, p. 335]), we see that $\phi_{N}(a_{1}\bmod N)\phi_{N}(a_{2}\bmod N)=\phi_{N}(a_{1}a_{2}\bmod N)$, so $\phi_{N}$ is also a group homomorphism, hence an isomorphism. ∎
6.3.2. Computations and examples
In view of Proposition 6.8 it is remarkably simple to compute the set of special classes $ax^{2}+xy+cy^{2}]$ for square discriminants.
Lemma 6.9.
Suppose that $D=N^{2}$ is an odd square discriminant and write $N=2k+1$, $k\geq 1$. Then
$\{[ax^{2}+xy+cy^{2}]$ $:1-4ac=D\}=\phi_{N}(\{2\delta\beta\,\mathrm{mod}\,N:\delta,\beta\in\mathbb{Z},\beta\mid k,\delta\mid k+1\}).$
Proof.
The quadratic form $ax^{2}+xy+cy^{2}$ has discriminant $(2k+1)^{2}$ if and only if ac=k(k+1)ac=-k(k+1). As above, by Gauss’s lemma there exist integers $\alpha,\beta,\gamma,\delta$ such that
$ax^{2}+xy+cy^{2}=(\alpha x+\gamma y)(\beta x+\delta y).$
Thus, $\alpha\beta=a$, $\gamma\delta=c$, and $\beta\gamma+\alpha\delta=1$. Therefore, $\alpha\beta\gamma\delta=-k(k+1)$ and, up to switching the factors, we can assume $\alpha\delta=k+1$ and $\beta\gamma=-k$. Let $g$ be the matrix $\left(\begin{smallmatrix}\delta&\beta\\ -\gamma&\alpha\end{smallmatrix}\right)$; note that $g\in\operatorname{SL}_{2}(\mathbb{Z})$. Then
(6.8) $\begin{split}g.(ax^{2}+xy+cy^{2})&=(\alpha(\delta x-\gamma y)+\gamma(\beta x+\alpha y))(\beta(\delta x-\gamma y)+\delta(\beta x+\alpha y))\\ &=x(2\delta\beta x+(2k+1)y)\\ &=Q_{N,2\delta\beta}\end{split}$
since $N=2k+1$. Conversely, for $\beta\mid k$ and $\delta\mid(k+1)$ the form $Q_{N,2\delta\beta}$ is equivalent to $\frac{k+1}{\delta}\beta x^{2}+xy-\frac{k}{\beta}\delta y^{2}$. ∎
Corollary 6.10.
If $N>3$ is an odd number, then there exists a knot with determinant $N^{2}$ having two Seifert surfaces that are not ambiently isotopic in $B^{4}$.
Proof.
Take $\beta=\delta=1$ in Lemma 6.9 so that the class $q]=\phi_{N}(2\bmod N)$ is of the form $ax^{2}+xy+cy^{2}]$ (as the proof of the lemma shows, that class is $(k+1)x^{2}+xy-ky^{2}]$). Since $\phi_{N}$ is injective, the class $q]^{2}=\phi_{N}(4\bmod N)$ is not the identity when $N>3$. We then apply Theorem 1.2 to obtain the desired result. ∎
7. Open Problems
7.1. Purely number-theoretic problems
As previously noted, analyzing the class group for positive discriminants is much harder than for negative discriminants (see e.g. Gauss’s class number problem). This is also illustrated in the strength of Corollary 1.3 (see also Lemma 6.1).
Problem 7.1.
Characterize the positive discriminants $D$ for which a class $ax^{2}+xy+cy^{2}]$ for $a,c\in\mathbb{Z}$, D=14acD=1-4ac, exists with $ax^{2}+xy+cy^{2}]^{2}\neq\mathrm{id}$.
If $D$ is a square, this is addressed in Corollary 6.10. For fundamental discriminants (i.e. $D$ square-free and $D\equiv 1\,\mathrm{mod}\,4$), the following problem is a weakening of the above.
Problem 7.2.
Show that there are infinitely many positive fundamental discriminants $D$, $D\equiv 1\,\mathrm{mod}\,4$, for which a class $ax^{2}+xy+cy^{2}]$ with $a,c\in\mathbb{Z}$, D=14acD=1-4ac, exists such that $ax^{2}+xy+cy^{2}]^{2}\neq\mathrm{id}$. If there are infinitely many such discriminants, establish or estimate the density of the set of such discriminants.
The problems addressed in this article are easily posed for higher genera as follows. Let $U$ be the bilinear form on $\mathbb{Z}^{4g}$ given by $U(e_{i},e_{j})=1$ if $(i,j)\in\{(1,2),(3,4),\ldots\}$ and 0 otherwise. Let $\theta$ (resp. $Q$) be the antisymmetrization (resp. symmetrization) of $U$. We call a rank $2g$ direct summand $L$ of $\mathbb{Z}^{4g}$ symplectic if there exists a basis $v_{1},\ldots,v_{2g}$ of $L$ such that for $i<j$ we have $\theta(v_{i},v_{j})=1$ if $(i,j)\in\{(1,2),(3,4),\ldots\}$ and 0 otherwise.
Problem 7.3.
For $g>1$, characterize the pairs $(q_{1},q_{2})$ of integral quadratic forms in $2g$ variables so that there exists a symplectic rank $2g$ direct summand $L$ of $\mathbb{Z}^{4g}$ such that $Q|_{L}$ is $\mathrm{Sp}_{2g}(\mathbb{Z})$-equivalent to $q_{1}$ and $Q|_{L^{\perp\kern-4.2pt\perp}}$ is $\mathrm{Sp}_{2g}(\mathbb{Z})$-equivalent to $q_{2}$. Here, as before $L^{\perp\kern-4.2pt\perp}=\{v\in\mathbb{Z}^{4g}:\theta(v,w)=0\text{ for all }w\in L\}$.
As phrased, we suspect Problem 7.3 to be difficult, though various weakenings of it appear amenable, for instance, to equidistribution methods as described in §1.4. For instance, one can ask for which discriminants a pair $(q_{1},q_{2})$ exists.
Example 7.4.
Pairs of forms $(q_{1},q_{2})$ as in Problem 7.3 may be constructed as follows: suppose that $(q_{1,i},q_{2,i})$ for $1\leq i\leq g$ are pairs of IBQFs so that $a_{i}x^{2}+xy+c_{i}y^{2}]^{2}*[q_{1,i}]=[q_{2,i}]$ for some $a_{i},c_{i}\in\mathbb{Z}$. By the present article, the pairs $(q_{1,i},q_{2,i})$ may be realized as in Problem 7.3 for $g=1$. It is then easy to see that one may realize the pair $(q_{1,1}\oplus\dots\oplus q_{1,g},q_{2,1}\oplus\dots\oplus q_{2,g})$. For a concrete example, one can set
(7.1) $q_{1}=\sum_{i=1}^{g}(x_{i}^{2}+x_{i}y_{i}+6y_{i}^{2}),\quad q_{2}=\sum_{i=1}^{g}(2x_{i}^{2}+x_{i}y_{i}+3y_{i}^{2}).$
Observe that $q_{1},q_{2}$ as in (7.1) are not in the same $\operatorname{GL}_{2g}(\mathbb{Z})$-class because $q_{2}$ does not represent $1$ (see also Example 6.2). We remark that a prototypical pair as in Problem 7.3 should not be of the special form considered in this example.
7.2. Topology questions
Livingston’s question [21] “Do there exist knots with two Seifert surfaces of the same genus $g$ that remain non-isotopic when pushed into the $4$-ball?” is resolved for all $g\geq 1$.
Example 7.5.
The example of $q_{1}$ and $q_{2}$ in (7.1) was pointed out to the authors by Kyle Hayden for its topological significance. Let $F_{1}$ and $F_{2}$ be any pair of genus $g$ Seifert surfaces for the same knot realizing $q_{1}$ and $q_{2}$ as their Seifert form. (Such exist by the $g\geq 1$ analogue of Lemma 5.3 or, to be concrete, one may take $F_{1}$ and $F_{2}$ as the $g$-fold boundary connected sum of $G_{1}$ and the $g$-fold boundary connected sum of $G_{2}$ respectively, where $G_{1}$ and $G_{2}$ are disjoint genus 1 pairs that have $x^{2}+xy+6y^{2}$ and $2x^{2}+xy+3y^{2}$ as their Seifert forms.) Then $\Sigma(F_{1}^{\mathbf{B}^{4}})$ and $\Sigma(F_{2}^{\mathbf{B}^{4}})$ have non-isometric intersection forms and, in particular, the pushed-in surfaces $F_{1}^{\mathbf{B}^{4}}$ and $F_{2}^{\mathbf{B}^{4}}$ are not topologically ambiently isotopic.
We wonder whether more subtle versions of Livingston’s question can also be answered affirmatively. For example, fixing $g\geq 1$ and a determinant $D\in\mathbb{Z}$, one may wonder, whether there exist genus $g$ Seifert surfaces $F_{1}$ and $F_{2}$ for the same knot with determinant $D$ such that $F_{1}^{\mathbf{B}^{4}}$ and $F_{2}^{\mathbf{B}^{4}}$ are not topologically ambiently isotopic. We pose a more refined question.
Problem 7.6.
Consider bilinear forms $U:\mathbb{Z}^{2g}\times\mathbb{Z}^{2g}\to\mathbb{Z}$ that antisymmetrize to a form of determinant $1$. For which such $U$ do there exist genus $g$ Seifert surfaces $F_{1}$ and $F_{2}$ for the same knot such that the Seifert form of $F_{1}$ is isometric to $U$ and the oriented $4$-manifolds $\Sigma(F_{1}^{\mathbf{B}^{4}})$ and $\Sigma(F_{2}^{\mathbf{B}^{4}})$ have non-isometric intersection form?
Problem 7.6 with the additional requirement of $F_{1}$ and $F_{2}$ being disjoint in their interior is equivalent to asking which quadratic forms $q$ arise as one of the entries of a pair as in Problem 7.3.
Let us return to genus $g=1$. For which pairs $s_{1}$ and $s_{2}$ of (equivalence classes of) IBQFs of the same discriminant $D\equiv 1\,\mathrm{mod}\,4$ is there a knot $K$ with two genus one Seifert surfaces $\Sigma_{1}$, $\Sigma_{2}$ having Seifert forms $s_{1}$, $s_{2}$ respectively? We have answered this question completely under the additional assumption that $\Sigma_{1}$ and $\Sigma_{2}$ are disjoint in Theorem 1.1. Without this restriction, we have the necessary condition that $\Sigma_{1},\Sigma_{2}$ must be S-equivalent, but we do not know if this is sufficient. In the context of this paper, S-equivalence of binary quadratic forms is quickest described by saying that $s_{1}$ and $s_{2}$ are S-equivalent if there exists an $s\in\mathcal{S}^{+}_{D}$ such that $s*s_{1}=s_{2}$. Hence, the relevant remaining question, which is equivalent to the conjecture stated in §5.4, is the following.
Problem 7.7.
Fix an integer $D$ such that $D\equiv 1\pmod{4}$, and let $s\in\mathcal{S}^{+}_{D}$ and $s_{1},s_{2}\in\mathcal{Q}^{+}_{D}$ be such that $s*s_{1}=s_{2}$. Do there exist genus $1$ Seifert surfaces $F_{1}$ and $F_{2}$ with boundary the same knot such that $s_{1}=s_{F_{1}}$ and $s_{2}=s_{F_{2}}$?
Given that in this text, we obstruct ambient isotopy of $F_{1}^{\mathbf{B}^{4}}$ and $F_{2}^{\mathbf{B}^{4}}$ in $\mathbf{B}^{4}$ using that $s_{F_{1}}\notin\{s_{F_{2}},\overline{s_{F_{2}}}\}$ are different, the following is a natural problem to consider.
Problem 7.8.
Do there exist genus $1$ Seifert surfaces $F_{1}$ and $F_{2}$ with boundary the same knot such that $s_{F_{1}}=s_{F_{2}}$ and $F_{1}^{\mathbf{B}^{4}}$ and $F_{2}^{\mathbf{B}^{4}}$ are not topologically ambiently isotopic in $\mathbf{B}^{4}$?